### v1

# Inverse problem for semidiscrete stochastic parabolic operators in arbitrary dimensions

Abstract.
In this paper, we study two types of inverse problems for space semidiscrete stochastic parabolic equations in arbitrary dimensions. The first problem concerns a semidiscrete inverse source problem, which involves determining the random source term of the white noise in the stochastic parabolic equation using observation data of the solution at the terminal time and the trace of the discrete spatial derivative of the solution on a subdomain at the boundary of the space. The second problem addresses a semidiscrete Cauchy inverse problem, which involves determining the solution of the stochastic parabolic equation in a subdomain of space-time, from the data of the solution and the trace of the discrete spatial derivative on a subdomain of the boundary of space and time. To address these problems, we establish two Carleman estimates for a semidiscrete stochastic parabolic operator in arbitrary dimensions: one for the homogeneous boundary condition and the other for the nonhomogeneous boundary condition. Applying these Carleman estimates, we obtain Lipschitz and Hölder stability for the first and second inverse problems, respectively.
Key words and phrases:
semidiscrete stochastic parabolic equations, global Carleman estimate, Inverse problems.
2020 Mathematics Subject Classification:
35R30, 35R60, 65C30, 60H15
1. Introduction
Let $(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P})$ be a complete filtered probability space on which a one-dimensional standard Brownian motion $\{B(t)\}_{t\geq 0}$ is defined such that $\{\mathcal{F}_{t}\}_{t\geq 0}$ is the natural filtration generated by $B(t)$, augmented by all $\mathbb{P}$ null sets in $\mathcal{F}$ and $\mathcal{X}$ a Banach space. The space $L^{2}_{\mathcal{F}}(\Omega;L^{2}(0,T;\mathcal{X}))$ consists of all $\mathcal{X}$-valued $\{\mathcal{F}_{t}\}_{t\geq 0}$-adapted processes $X(\cdot)$ such that $\mathbb{E}\left(\int_{0}^{T}\|X(\cdot)\|^{2}_{\mathcal{X}}\,dt\right)<\infty.$ The space $L^{2}_{\mathcal{F}}(\Omega;L^{\infty}(0,T;\mathcal{X}))$ consists of all $\mathcal{X}$-valued $\{\mathcal{F}_{t}\}_{t\geq 0}$-adapted bounded processes. Finally, $L^{2}_{\mathcal{F}}(\Omega;C([0,T];\mathcal{X}))$ denotes the space of all $L^{p}_{h}(\mathcal{M})$-valued $\{\mathcal{F}_{t}\}_{t\geq 0}$-adapted continuous processes $X(\cdot)$ such that $\mathbb{E}\left(\sup_{t\in[0,T]}\|X(\cdot)\|^{2}_{\mathcal{X}}\right)<\infty.$ For simplicity, we often write $L^{2}_{\mathcal{F}}(0,T;\mathcal{X})$ or $L^{\infty}_{\mathcal{F}}(0,T;\mathcal{X})$ whenever no confusion may arise.
Let $T>0$ and $n\in\mathds{N}$, and $G\triangleq(0,1)^{n}$. We consider the following stochastic parabolic equation:
[…]
In (1.1), $y$ denotes the state variable, $g$ is a source term representing the intensity of the random force of white noise, and $\gamma^{i},a^{1i}$, and $a^{k}$ (for $i=1,\ldots,n$ and $k=2,3$) are given functions. We refer to KR77 , Zho92 , and lu2021mathematical for the existence and uniqueness results of (1.1).
The main objective of this paper is to study the inverse random source problem and the Cauchy problem associated with the semidiscrete spatial approximation of the stochastic parabolic equation (1.1). To this end, we introduce the spatial discretization employed throughout this work.
Given a set of points $\mathcal{M}\subseteq G\cap h\mathbb{Z}^{n}$, called the primal mesh, we define the dual mesh in the direction $e_{k}$, where $\{e_{k}\}_{k=1}^{n}$ denotes the standard basis of $\mathbb{R}^{n}$, by
$\mathcal{M}_{k}^{\ast}\triangleq\left\{x\pm\frac{h}{2}e_{k}\mid x\in\mathcal{M}\right\}.$
This allows us to define the boundary points in the $e_{k}$ direction as $\partial_{k}\mathcal{M}\triangleq(\mathcal{M}^{\ast}_{k})^{\ast}_{k}\setminus\mathcal{M}$, and we set
$\partial\mathcal{M}\triangleq\bigcup_{k=1}^{n}\partial_{k}\mathcal{M},\quad\textrm{ and }\quad\overline{\mathcal{M}}\triangleq\bigcup_{k=1}^{n}(\mathcal{M}^{\ast}_{k})^{\ast}_{k}.$
We denote by $C(\mathcal{M})$ the set of real-valued functions defined in the mesh $\mathcal{M}$.
The outward normal to $\mathcal{M}$ in the direction $e_{i}$ is defined as $\nu_{i}\in C(\partial\mathcal{M}_{i})$:
(1.2) $\forall x\in\partial_{i}\mathcal{M},\quad\nu_{i}(x)\triangleq\begin{cases}1&\mbox{if }x-\tfrac{h}{2}e_{i}\in\mathcal{M}_{i}^{\ast}\ \mbox{and }x+\tfrac{h}{2}e_{i}\notin\mathcal{M}_{i}^{\ast},\\ -1&\mbox{if }x-\tfrac{h}{2}e_{i}\notin\mathcal{M}_{i}^{\ast}\ \mbox{and }x+\tfrac{h}{2}e_{i}\in\mathcal{M}_{i}^{\ast},\\ 0&\mbox{otherwise}.\end{cases}$
We also define the trace operator $t_{r}^{i}$ for $u\in C(\mathcal{M}^{\ast}_{i})$ by
(1.3) $\forall x\in\partial_{i}\mathcal{M},\quad t_{r}^{i}(u)(x)\triangleq\begin{cases}u\!\left(x-\tfrac{h}{2}e_{i}\right)&\nu_{i}(x)=1,\\ u\!\left(x+\tfrac{h}{2}e_{i}\right)&\nu_{i}(x)=-1,\\ 0&\nu_{i}(x)=0.\end{cases}$
Finally, to obtain the discrete version of our system via the finite difference method, we introduce the difference and average operators in the direction $e_{k}$:
$(D_{k}u)(x)\triangleq\frac{\tau_{+k}u(x)-\tau_{-k}u(x)}{h},\quad(A_{k}u)(x)\triangleq\frac{\tau_{+k}u(x)+\tau_{-k}u(x)}{2},\quad\text{for }x\in\mathcal{M}^{\ast}_{k},$
where $\tau_{\pm k}u(x)\triangleq u\!\left(x\pm\frac{h}{2}e_{k}\right)$.
The Banach spaces $\mathcal{X}$ required for the formulation of both inverse problems are defined as follows. We denote by $L^{p}_{h}(\mathcal{M})$ with $1\leq p<\infty$ the space of function defined on $\mathcal{M}$ with the associated norm.
$\|u\|_{L_{h}^{p}(\mathcal{M})}^{p}\triangleq\int_{\mathcal{M}}|u|^{p},$
where the integral is defined by
$\int_{\mathcal{M}}|u|^{p}\triangleq h^{n}\sum_{x\in\mathcal{M}}|u(x)|^{p}.$
Analogously, we denote by $W^{1,p}_{h}(\mathcal{M})$, $W_{h}^{2,p}(\mathcal{M})$, and $L^{\infty}_{h}(\mathcal{M})$ the set of functions defined in $\mathcal{M}$ with the associated norms.
$\|u\|_{W_{h}^{1,p}(\mathcal{M})}^{p}\triangleq\|u\|_{L_{h}^{p}(\mathcal{M})}^{p}+\sum_{i\in\llbracket 1,n\rrbracket}\int_{\mathcal{M}_{i}^{\ast}}|D_{i}u|^{p},$
$\|u\|_{W_{h}^{2,p}(\mathcal{M})}^{p}\triangleq\|u\|_{W_{h}^{1,p}(\mathcal{M})}^{p}+\sum_{i,j\in\llbracket 1,n\rrbracket}\int_{Q_{ij}^{\ast}}|D_{ij}^{2}u|^{p},$
and
$\|u\|_{L_{h}^{\infty}(\mathcal{M})}\triangleq\max_{x\in\mathcal{M}}|u(x)|,$
respectively. In particular, for $p=2$, we adopt the usual notation $W^{1,2}_{h}(\mathcal{M})=H_{h}^{1}(\mathcal{M})$ and $W_{h}^{2,2}(\mathcal{M})=H_{h}^{2}(\mathcal{M})$.
For the boundary, we proceed as above, that is, for $u\in C(\partial\mathcal{M})$ we define the integral over the boundary $\partial\mathcal{M}$ as
$\int_{\partial\mathcal{M}}u\triangleq\sum_{i=1}^{n}\int_{\partial_{i}\mathcal{M}}u(x),$
where
$\int_{\partial_{i}\mathcal{M}}u(x)\triangleq h^{n-1}\sum_{x\in\partial_{i}\mathcal{M}}u(x).$
Then, we can denote as $L^{2}_{h}(\partial\mathcal{M})$ and $H^{1}_{h}(\partial\mathcal{M})$ the set of functions defined on $\partial\mathcal{M}$ with the associated norm
$\left\|u\right\|^{2}_{L^{2}_{h}(\partial\mathcal{M})}\triangleq\int_{\partial\mathcal{M}}|u|^{2},$
and
$\left\|u\right\|^{2}_{H^{1}_{h}(\partial\mathcal{M})}\triangleq\sum_{i\in\llbracket 1,n\rrbracket}\left(\int_{\partial_{i}\mathcal{M}}|u|^{2}+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\int_{(\partial_{i}\mathcal{M})_{j}^{\ast}}|D_{j}u|^{2}\right),$
respectively. Finally, following ervedoza-2011 , we adopt the discrete $H^{1/2}_{h}$-norm defined for $v\in C(\partial\mathcal{M})$ by
$\|v\|_{H_{h}^{1/2}(\partial\mathcal{M})}\triangleq\min_{\left.u\right|_{\partial\mathcal{M}}=v}\|u\|_{H^{1}_{h}(\mathcal{M})}.$
Following the same reasoning, we analogously define the semidiscrete spaces $L^{2}_{\mathcal{F}}(0,T;L_{h}^{\infty}(\mathcal{M}))$, $L^{2}_{\mathcal{F}}(0,T;H_{h}^{1}(\mathcal{M}))$, $L^{2}_{\mathcal{F}}(0,T;L_{h}^{2}(\partial\mathcal{M}))$, $L^{2}_{\mathcal{F}}(0,T;H_{h}^{1}(\partial\mathcal{M}))$, and $L^{2}_{\mathcal{F}}(0,T;H_{h}^{1/2}(\partial\mathcal{M}))$.
Given the previous definitions, we are now ready to formulate the problems to be addressed. Let $\gamma_{i}$ denote the sampling of $\gamma^{i}$ onto both the primal and dual meshes. Furthermore, let $a_{1i}$, $a_{2}$, $a_{3}$, and $a_{4}$ represent the evaluations of $a^{1i}$, $a^{2}$, and $a^{3}$ on the primal mesh, for $i=1,\ldots,n$, respectively. We now formulate the discrete versions of the inverse problems considered in this paper, namely, the discrete inverse random source problem and the discrete Cauchy problem for the spatially semidiscrete stochastic parabolic equation in $n$ dimensions.
Semidiscrete inverse random source problem. We consider the following $n$-dimensional semidiscrete stochastic parabolic equation, derived from (1.1), assuming $a_{3}=0$:
[…]
where the second-order operator is defined by $\mathcal{A}_{h}w\triangleq dw-\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}(w))dt$, and $w_{0}$ is the initial condition.
We define the observation operator $\Lambda_{1}:L^{2}_{\mathcal{F}}(0,T;H_{h}^{1}(\mathcal{M}))\longrightarrow X_{1}$ as
$\Lambda_{1}(g)\triangleq\left(\left.\partial_{\nu}w\right|_{\Gamma^{+}\times(0,T)},w(T)\right),$
where the normal derivative is given by
$\partial_{\nu}w\triangleq\sum_{i=1}^{n}t_{r}^{i}(D_{i}w)e_{i}\nu_{i}.$
The set $\Gamma^{+}\subset\partial\mathcal{M}$ is defined as
$\Gamma^{+}\triangleq\left\{x\in\partial\mathcal{M}:x\in\partial_{i}\mathcal{M}\text{ and }(x-x^{*})e_{i}\nu_{i}\geq 0\right\},$
for some fixed point $x^{*}\notin\overline{\mathcal{M}}$, and the norm that will be adopted in $X_{1}\triangleq\left(L^{2}_{\mathcal{F}}(0,T;L_{h}^{2}(\Gamma^{+}))\right)^{n}\times L^{2}_{\mathcal{F}_{T}}(\Omega;L_{h}^{2}(\mathcal{M}))$ is
$\left\|\left(f_{1},...,f_{n},f_{n+1}\right)\right\|_{X_{1}}\triangleq\sum_{i=1}^{n}\|f_{i}\|_{L^{2}_{\mathcal{F}}(0,T;L^{2}_{h}(\Gamma^{+}))}+\|f_{n+1}\|_{L^{2}_{\mathcal{F}_{T}}(\Omega;L^{2}_{h}(\mathcal{M}))},$
for $f_{i}\in L^{2}_{\mathcal{F}}(0,T;L^{2}_{h}(\mathcal{M}))$ with $i=1\cdots n$ and $f_{n+1}\in L^{2}_{\mathcal{F}_{T}}(\Omega;L^{2}_{h}(\mathcal{M}))$.
The direct problem consists of solving (1.4) for a given $g$ to compute $\Lambda_{1}(g)$. The inverse problem is to recover $g$ from the knowledge of $\Lambda_{1}(g)$.
We present a conditional stability result.
Theorem 1.1.
Let $a_{1i}\in L^{\infty}_{\mathcal{F}}(0,T;L_{h}^{\infty}(\mathcal{M}))$, $a_{2}\in L_{\mathcal{F}}^{\infty}(0,T;L_{h}^{n^{\ast}}(\mathcal{M}))$ where the exponent $n^{\ast}$ satisfies
(1.5) $\begin{cases}n^{\ast}>2,&\quad n=2,\\ n^{\ast}\geq n,&\quad n\geq 3.\end{cases}$
and assume that $g_{k}\in L^{2}_{\mathcal{F}}(0,T;H_{h}^{1}(\mathcal{M}))$ for $k=1,2$, which satisfy
(1.6) $|D_{i}(g_{1}-g_{2})|\leq C|A_{i}(g_{1}-g_{2})|,\quad\text{in }Q_{i}^{*},\;\mathbb{P}\text{-a.s.},\quad\text{for }i=1,\dots,n.$
Then, there exists a constant $C>0$, independent of $h$, such that
(1.7) $\|g_{1}-g_{2}\|_{L^{2}_{\mathcal{F}}(0,T;L_{h}^{2}(\mathcal{M}))}\leq C\|\Lambda_{1}(g_{1})-\Lambda_{1}(g_{2})\|_{X_{1}}.$
Remark 1.2.
The uniqueness follows directly from Theorem 1.1: if $\Lambda_{1}(g_{1})=\Lambda_{1}(g_{2})$ $\mathbb{P}$-a.s., then $g_{1}=g_{2}$ in $Q$, $\mathbb{P}$-a.s.
Semidiscrete Cauchy Problem. We now consider the same semi-discretization of (1.1), but with $g=0$ and a non-homogeneous Dirichlet boundary condition:
[…]
Although the discretization is structurally similar to the previous problem, the key distinction lies in the type of data: Here, we observe the boundary values and aim to recover the interior solution.
We define the observation operator $\Lambda_{2}:L^{2}_{\mathcal{F}}(0,T;H_{h}^{2}(\mathcal{M}))\to X_{2}$ by
$\left.\Lambda_{2}(w)\triangleq\left(w,\partial_{\nu}w\right)\right|_{\Gamma^{+}\times(0,T)},$
where $w$ solves (1.8). In this case, for the space $X_{2}\triangleq H^{1}_{\mathcal{F}}(0,T;H_{h}^{1}(\partial\mathcal{M})\cap H_{h}^{1/2}(\partial\mathcal{M}))\times\left(L^{2}_{\mathcal{F}}(0,T;L_{h}^{2}(\Gamma^{+}))\right)^{n}$ adopted the norm defined for
$\left\|\left(f_{1},f_{2},...,f_{n},f_{n+1}\right)\right\|_{X_{2}}\triangleq\|f_{1}\|_{H^{1}_{\mathcal{F}}(0,T;H_{h}^{1}(\partial\mathcal{M})\cap H_{h}^{1/2}(\partial\mathcal{M}))}+\sum_{i=2}^{n+1}\|f_{i}\|_{L^{2}_{\mathcal{F}}(0,T;L_{h}^{2}(\Gamma^{+}))},$
for $f_{1}\in H^{1}_{\mathcal{F}}(0,T;H_{h}^{1}(\partial\mathcal{M})\cap H_{h}^{1/2}(\partial\mathcal{M}))$ and $f_{i}\in L^{2}_{\mathcal{F}}(0,T;L_{h}^{2}(\Gamma^{+}))$ with $i=2\cdots n+1$.
The inverse problem consists of determining the solution $w$ in a subdomain $\mathcal{M}_{0}\subset\mathcal{M}$ over the time interval $(\varepsilon,T-\varepsilon)$, based on the boundary measurement given by $\Lambda_{2}(w)$. We establish the following Hölder-type stability estimate.
Theorem 1.3.
Let $a_{1i},a_{2}\in L^{\infty}_{\mathcal{F}}(0,T;L_{h}^{\infty}(\mathcal{M}))$, $a_{3}\in L^{\infty}_{\mathcal{F}}(0,T;W_{h}^{1,\infty}(\mathcal{M}))$, $\xi\in H^{1}_{\mathcal{F}}(0,T;H_{h}^{1}(\partial\mathcal{M})\cap H_{h}^{1/2}(\partial\mathcal{M)})$, and let $\varepsilon>0$. Then there exist constants $C>0$, $h^{*}>0$, and $\kappa\in(0,1)$ such that for all $h\in(0,h^{*})$ and any solution $w\in L^{2}_{\mathcal{F}}(0,T;H_{h}^{2}(\mathcal{M}))$ of (1.8) satisfying
$\|w\|_{L^{2}_{\mathcal{F}}(0,T;H_{h}^{2}(\mathcal{M}))}\leq\mathbf{M},$
the following estimate holds
(1.9) $\|w\|_{L^{2}_{\mathcal{F}}(\varepsilon,T-\varepsilon;H_{h}^{2}(\mathcal{M}_{0}))}\leq C\max{\left\{\|\Lambda_{2}(w)\|_{X_{2}},\mathbf{M}e^{-Ch^{-1}},\mathbf{M}^{\kappa}\|\Lambda_{2}(w)\|_{X_{2}}^{1-\kappa}\right\}}.$
Remark 1.4.
The size of the observation set $\Gamma^{+}$ depends on the choice of the reference point $x^{*}$. If $x^{*}$ is chosen so that the vector $x-x^{*}$ is perpendicular to the face of $\partial\mathcal{M}$, then $\Gamma^{+}$ is smaller. In contrast, if $x^{*}$ is chosen in a more oblique direction, $\Gamma^{+}$ increases in size. Thus, an appropriate choice of $x^{*}$ could minimize the required observation region.
Remark 1.5.
We observe that if we consider $h$ to be small enough and the norm $|\Lambda_{2}(w)|{X_{2}}$ to be small compared to the constant $\mathbf{M}$ (i.e., $|\Lambda_{2}(w)|{X_{2}}<\mathbf{M}$), we can recover the estimate presented in WWW:2024 . In that reference, the first two terms in (1.9) were neglected.
The two inverse problems under consideration are associated with stochastic parabolic equations. Stochastic partial differential equations (SPDEs) have broad applications across multiple disciplines. They model phenomena such as particle diffusion and heat propagation in physics, option pricing in finance, epidemic spread and population dynamics in biology, systems with uncertain parameters in engineering, and pollutant dispersion in environmental sciences. These equations incorporate randomness and uncertainty, making them particularly suitable for modeling inherently unpredictable phenomena. Consequently, stochastic PDEs often provide more realistic representations than their deterministic counterparts. Nevertheless, research on inverse problems for stochastic PDEs remains limited. For further details, see (lu2021mathematical, , Chapter 5) and the references therein.
For the inverse source problem related to stochastic parabolic equations, ref27 establishes uniqueness in determining the source function $f$ in the drift term. The inverse source problem of simultaneously determining two types of sources, $f$ (drift term) and $g$ (diffusion term), in a stochastic parabolic equation was studied in ref37 using observations at the final time and on the lateral boundary. A unique result for this inverse problem was obtained using a global Carleman estimate for the continuous stochastic parabolic equation. In addition, ref1 and ref2 discuss the application of regularization techniques within numerical methods to inverse random source problems.
The Cauchy problem aims to recover the solution using the observed data on the lateral boundary. In Yamamoto_2009 , conditional stability was established for the Cauchy problem of deterministic parabolic equations. The conditional stability and the convergence rate of the Tikhonov regularization method for the Cauchy problem of stochastic parabolic equations were obtained in ref13 . It is worth mentioning that ref28 provides a detailed review of inverse problems for stochastic PDEs. However, all of these results were obtained within a continuous framework. For semidiscrete inverse problems related to stochastic differential equations, see WWW:2024 , which studies the one-dimensional case of both issues, including a semidiscrete inverse random source problem and a semidiscrete Cauchy problem.
This article builds upon the results of WWW:2024 , extending them to arbitrary spatial dimensions and providing several refinements. In particular, Theorem 1.1 provides integrability conditions for the coefficient $a_{2}$ that depend on the spatial dimension of the problem. Furthermore, Theorem 1.3 establishes a stability estimate in which the right-hand side of inequality (1.9) explicitly reveals the dependence on the mesh size $h$ and the size of the observation operator. As a consequence, the second result recovers the same estimate as in WWW:2024 when $h$ and the observation region are chosen sufficiently small.
The main tools to address the previous inverse problems are two Carleman estimates. The first is for a homogeneous Dirichlet condition.
Theorem 1.6.
Let $\mbox{reg}^{0}>0$, $\psi$ satisfy assumptions (2.1) and $\varphi$ according to (2.2). For $\lambda_{1}\geq 1$ sufficiently large, there exist $C$, $\tau_{0}\geq 1$, $h_{0}>0$, $\varepsilon_{0}>0$, depending on $G$, $\beta$, $\mbox{reg}^{0}$, $T$, and $\lambda$, such that for any $\gamma$, with $\mbox{reg}(\gamma)\leq\mbox{reg}^{0}$ we have
(1.10) $\begin{split}&J(w)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|w|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}w|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|g|^{2}dt\\ &\leq\,C_{s_{0},\varepsilon}\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}|f|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|D_{i}g|^{2}dt+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}st_{r}^{i}(e^{2s\varphi}|D_{i}w|^{2})\,dt\right.\\ &\left.+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|w|^{2}\right|_{t=T}\right).\end{split}$
for all $\tau\geq\tau_{0}$, $0<h\leq h_{0}$, $s(t)h\leq\epsilon_{0}$, where $J(w)\triangleq\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}s\lambda_{1}^{2}\varphi e^{2s\theta\varphi}\,|D_{i}w|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi e^{2s\theta\varphi}|A_{i}D_{i}w|^{2}\,dt$, $f\in L^{2}_{\mathcal{F}}(0,T;L^{2}(\mathcal{M}))$, $g\in L^{2}_{\mathcal{F}}(0,T;H^{1}(\mathcal{M}))$, and $w$ satisfies $dw-\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}w)\,dt=fdt+gdB(t)$ with $w=0$ on $\partial\mathcal{M}$ and $w(0)=0$ in $\mathcal{M}$.
The second is for the nonhomogeneous Dirichlet condition given in Theorem 3.3. Theorem 1.6 and Theorem 3.3 extend the one-dimensional Carleman estimates presented in WWW:2024 . We adopt the weight function introduced in that work. However, in the multidimensional setting, additional difficulties arise due to the need to incorporate second-order terms $D^{2}_{ij}$ and the fact that the boundary is no longer a simple pair of points. To establish these estimates, we rely on fundamental discrete tools, such as integration by parts and the product rule, which we state below.
Proposition 1.7 ((LOPD, , Lemma 2.2)).
For any $v\in C(\mathcal{M}_{i}^{\ast})$ and $u\in C(\overline{\mathcal{M}}_{i})$, the following identities hold: for the difference operator
(1.11) $\int_{\mathcal{M}}u\,D_{i}v=-\int_{\mathcal{M}_{i}^{\ast}}v\,D_{i}u+\int_{\partial_{i}\mathcal{M}}u\,t_{r}^{i}(v)\nu_{i},$
and for the average operator
(1.12) $\int_{\mathcal{M}}u\,A_{i}v=\int_{\mathcal{M}_{i}^{\ast}}v\,A_{i}u-\frac{h}{2}\int_{\partial_{i}\mathcal{M}}u\,t_{r}^{i}(v).$
Next, we recall the discrete product rule, which relates the average and difference operators.
Proposition 1.8 ((boyer-2010-elliptic, , Lemma 2.1 and 2.2)).
For $u,v\in C(\overline{\mathcal{M}})$, the following identities hold in $\mathcal{M}^{\ast}_{i}$. For the difference operator
(1.13) $D_{i}(u\,v)=D_{i}u\,A_{i}v+A_{i}u\,D_{i}v,$
and for the average operator
(1.14) $A_{i}(u\,v)=A_{i}u\,A_{i}v+\frac{h^{2}}{4}D_{i}u\,D_{i}v.$
Moreover, on $\mathcal{M}$} we have
(1.15) $u=A^{2}_{i}u-\frac{h^{2}}{4}D^{2}_{i}u.$
The remainder of this paper is organized as follows. In Section 2, we prove a Lipschitz stability result of the space-semidiscrete inverse source problem as a direct result of the Carleman estimate presented in the Appendix B. Finally, in Section 3, we study the space-semidiscrete inverse Cauchy problem and establish a Hölder stability result based on a second Carleman estimate and a geometric argument.
2. Semidiscrete inverse random source problem
In this section, we establish the stability of the discrete inverse random source problem (Theorem 1.1). The key tool for proving this result is a semidiscrete Carleman estimate, whose proof is deferred to B. To formulate this estimate, we first introduce the necessary assumptions on $\gamma\triangleq(\gamma_{1},...,\gamma_{n})$ and the associated weight functions. Assume $\gamma_{i}>0$ for all i=1,,ni=1,...,n with
$\mbox{reg}(\gamma)\triangleq\text{ess}\sup_{\begin{subarray}{c}x\in G\\ i=1,\cdots,n\end{subarray}}\left(\gamma_{i}+\frac{1}{\gamma_{i}}+\sum_{i=1}^{n}|\partial_{i}\gamma_{i}|^{2}\right)<\infty.$
Now, let $x^{\ast}\notin\overline{G}$ and, for $\lambda>1$, define the weight functions
(2.1) $\psi(x)$ $=|x-x^{*}|^{2},$
(2.2) $\varphi(t,x)$ $=e^{\lambda\psi(x)},$
and for $t_{0}\in(0,T)$ and $\beta>0$,
(2.3) $\theta(t)=\beta|t-t_{0}|^{2}.$
Given $\tau\geq 1$, we set $s(t)=\tau e^{-\lambda\theta(t)}.$
Remark 2.1.
Compared to EstimateCarlemanStochasticControl and other works on parabolic operators, the weight function here has been modified, particularly in its temporal component; however, the structure $s(t)\varphi(x)$ is preserved. Moreover, the functions satisfy $|\theta(t)|\leq C$ and $|\theta_{t}(t)|\leq C$ for some constant $C>0$ and for all $t\in[0,T]$. As a result, the estimates in Appendix A are obtained, and most of the cross-product terms can be treated in a manner analogous to EstimateCarlemanStochasticControl , slightly simplifying the computations for the new Carleman estimate in this article.
2.1. Proof of Theorem 1.1.
Let $w_{1}$ and $w_{2}$ be the solutions of (1.4) with the random forcing term $g_{1}$ and $g_{2}$, respectively. Setting $\tilde{w}=w_{1}-w_{2}$ and $\tilde{g}=g_{1}-g_{2}$, we obtain
$\begin{cases}d\tilde{w}-\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}\tilde{w})=\left(\sum_{i\in\llbracket 1,n\rrbracket}a_{1i}A_{i}D_{i}(\tilde{w})+a_{2}\tilde{w}\right)\,dt+\tilde{g}\,dB(t),&\text{in}\,\mathcal{M}\times(0,T),\\ \tilde{w}=0,&\text{on}\,\partial\mathcal{M}\times(0,T),\\ \tilde{w}(0)=0,\quad&\text{in}\,\mathcal{M}.\end{cases}$
Applying Theorem 1.6 to $\tilde{w}$, we derive the following estimate.
$\begin{split}&J(\tilde{w})+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|\tilde{w}|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}w|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|\tilde{g}|^{2}dt\\ &\leq\,C_{s_{0},\varepsilon}\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}\left|\sum_{i=1}^{n}a_{1i}A_{i}D_{i}\tilde{w}+a_{2}\tilde{w}\right|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|D_{i}\tilde{g}|^{2}dt\right.\\ &\left.+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}st_{r}^{i}(e^{2s\varphi}|D_{i}\tilde{w}|^{2})\,dt+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|\tilde{w}|^{2}\right|_{t=T}\right).\end{split}$
where $J(w)$ is defined as $\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi e^{2s\theta\varphi}\,|D_{i}w|^{2}\,dt+\mathbb{E}\int_{Q}s\varphi e^{2s\theta\varphi}|A_{i}D_{i}w|^{2}\,dt$. By applying Young’s inequality for the first integral and using (1.6) for the second term on the right-hand side of the previous inequality, we can rewrite the inequality above as
(2.4) $\begin{split}&J(\tilde{w})+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|\tilde{w}|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}w|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|\tilde{g}|^{2}\,dt\\ &\leq C_{s_{0},\varepsilon}\left(\sum_{i=1}^{n}\mathbb{E}\int_{Q}\,e^{2s\varphi}|a_{1i}A_{i}D_{i}(\tilde{w})|^{2}\,dt+\mathbb{E}\int_{Q}\,e^{2s\varphi}|a_{2}\tilde{w}|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|A_{i}\tilde{g}|^{2}\,dt\right.\\ &\left.+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}st_{r}^{i}(e^{2s\varphi}|D_{i}\tilde{w}|^{2})\,dt+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|\tilde{w}|^{2}\right|_{t=T}\right).\end{split}$
Observe that, by applying Hölder’s inequality with $\frac{1}{r}+\frac{1}{q}=1$ to the second term on the right-hand side, we obtain
$\mathbb{E}\int_{Q}e^{2s\varphi}|a_{2}\tilde{w}|^{2}\,dt\leq\mathbb{E}\int_{0}^{T}\left(\int_{\mathcal{M}}|a_{2}|^{2r}\,\right)^{1/r}\left(\int_{\mathcal{M}}|e^{s\varphi}\tilde{w}|^{2q}\right)^{1/q}dt.$
Applying the Sobolev-type inequality stated in A.6 with $p=2$ and $p^{\ast}=2q$, we obtain the following.
$\mathbb{E}\int_{Q}e^{2s\varphi}|a_{2}\tilde{w}|^{2}\,dt\leq\mathbb{E}\int_{0}^{T}\left(\int_{\mathcal{M}}|a_{2}|^{2r}\,\right)^{2/r}\left(\int_{\mathcal{M}}|e^{s\varphi}\tilde{w}|^{2}+\sum_{i=1}^{n}\int_{\mathcal{M}}|D_{i}(e^{s\varphi}\tilde{w})|^{2}\right)dt,$
where we can see that choosing $n^{\ast}=2r$ satisfies condition (2.6). Taking into account (1.13) and the triangle inequality, from the above inequality, we deduce
$\mathbb{E}\int_{Q}e^{2s\varphi}|a_{2}\tilde{w}|^{2}\,dt\leq C\|a_{2}\|_{L^{\infty}_{\mathcal{F}}(0,T;L_{h}^{n^{\ast}}(\mathcal{M}))}$ $\mathbb{E}\int_{0}^{T}\left(\int_{\mathcal{M}}|e^{s\varphi}\tilde{w}|^{2}\right$.
$\left.+\sum_{i=1}^{n}\int_{\mathcal{M}_{i}^{\ast}}|D_{i}e^{s\varphi}|^{2}|A_{i}\tilde{w}|^{2}+|A_{i}e^{s\varphi}|^{2}|D_{i}\tilde{w}|^{2}\right)dt.$
Notice that thanks to (1.14), for $i=1,..,n$ and $u\in C(\mathcal{M})$ we can obtain the following inequality.
(2.5) $A_{i}(|u|^{2})\geq|A_{i}u|^{2}.$
Thus, from the above inequality, (1.12), and using the boundary condition $\tilde{w}=0$ on $\partial\mathcal{M}$, it follows that
$\int_{\mathcal{M}_{i}^{\ast}}|D_{i}e^{s\varphi}|^{2}|A_{i}\tilde{w}|^{2}\leq\int_{\mathcal{M}_{i}^{\ast}}|D_{i}e^{s\varphi}|^{2}A_{i}(|\tilde{w}|^{2})\leq\int_{\mathcal{M}}A_{i}(|D_{i}e^{s\varphi}|^{2})|\tilde{w}|^{2}$
Therefore,
$\mathbb{E}\int_{Q}e^{2s\varphi}|a_{2}\tilde{w}|^{2}\,dt\leq C\|a_{2}\|_{L^{\infty}_{\mathcal{F}}(0,T;L_{h}^{n^{\ast}}(\mathcal{M}))}$ $\mathbb{E}\int_{0}^{T}\left(\int_{\mathcal{M}}\left(e^{2s\varphi}+\sum_{i=1}^{n}A_{i}(|D_{i}e^{s\varphi}|^{2})\right)|\tilde{w}|^{2}\right$.
$\left.+\sum_{i=1}^{n}\int_{\mathcal{M}_{i}^{\ast}}|A_{i}e^{s\varphi}|^{2}|D_{i}\tilde{w}|^{2}\right)dt.$
Finally, using the bounds $A_{i}(|D_{i}(e^{s\varphi})|^{2})\leq\lambda^{2}s^{2}\varphi^{2}(\partial_{i}\psi)^{2}e^{2s\varphi}+e^{2s\varphi}\mathcal{O}_{\lambda}((sh)^{2})$ and $|A_{i}(e^{s\varphi})|^{2}\leq e^{2s\varphi}+e^{2s\varphi}\mathcal{O}_{\lambda}((sh)^{2})$, we conclude that
(2.6) $\begin{split}&\mathbb{E}\int_{Q}e^{2s\varphi}|a_{2}\tilde{w}|^{2}\,dt\leq C\|a_{2}\|_{L^{\infty}_{\mathcal{F}}(0,T;L_{h}^{n^{\ast}}(\mathcal{M}))}\mathbb{E}\int_{0}^{T}\left(\int_{\mathcal{M}}\left(1+\lambda^{2}s^{2}\varphi^{2}|\nabla\psi|^{2}\right)e^{2s\varphi}|\tilde{w}|^{2}\right.\\ &\left.+\sum_{i=1}^{n}\int_{\mathcal{M}_{i}^{\ast}}e^{2s\varphi}|D_{i}\tilde{w}|^{2}\right)dt+\mathcal{O}_{\lambda}((sh)^{2})\mathbb{E}\int_{0}^{T}\left(\int_{\mathcal{M}}e^{2s\varphi}|\tilde{w}|^{2}+\sum_{i=1}^{n}\int_{\mathcal{M}_{i}^{\ast}}e^{2s\varphi}|D_{i}\tilde{w}|^{2}\right)\,dt.\end{split}$
On the other hand, applying (2.5) and performing integration by parts with respect to the averaging operator (1.12), we obtain the following.
$\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|A_{i}\tilde{g}|^{2}\,dt\leq\mathbb{E}\int_{Q_{i}^{\ast}}e^{2s\varphi}A_{i}(|g|^{2})\,dt=\mathbb{E}\int_{Q}A_{i}(e^{2s\varphi})|\tilde{g}|^{2}\,dt+\frac{h}{2}\mathbb{E}\int_{\partial_{i}Q}t_{r}^{i}(e^{2s\varphi})|\tilde{g}|^{2}\,dt.$
Now, considering $\mathbb{E}\int_{\partial_{i}Q}|\tilde{g}|^{2}\,dt=\mathbb{E}\int_{\partial_{i}Q}|d\tilde{w}|^{2}=0$ and $e^{-2s\varphi}A_{i}(e^{2s\varphi})=1+\mathcal{O}((sh)^{2})$, where we used $\tilde{w}=0$ in $\partial_{i}Q$ and Proposition 2.4 in AA:perez:2024 , respectively, we obtain the following.
$\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|A_{i}\tilde{g}|^{2}\,dt\leq\mathbb{E}\int_{Q}\left(e^{2s\varphi}+\mathcal{O}_{\lambda}((sh)^{2})e^{2s\varphi}\right)|\tilde{g}|^{2}\,dt.$
Finally, recording that $s(t)\triangleq\tau\theta(t)$, using the above inequality in the third term on the right-hand side of (2.4) and letting $\tau\geq C$, we can absorb the initial three terms on the right-hand side of the same inequality (2.4). Therefore, we can achieve the following inequality:
$\begin{split}&J(\tilde{w})+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|\tilde{w}|^{2}\,dt+\sum_{i,j\in\llbracket 1,n\rrbracket}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}w|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|\tilde{g}|^{2}\,dt\\ &\leq C\left(\sum_{i\in\llbracket 1,n\rrbracket}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}st_{r}^{i}(e^{2s\varphi}|D_{i}\tilde{w}|^{2})\,dt+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|\tilde{w}|^{2}\right|_{t=T}\right).\end{split}$
From the above, we deduce (1.7) and the proof of the theorem is completed.
3. Semidiscrete Cauchy problem.
In this section, we establish the conditional stability result for the semidiscrete Cauchy problem, as stated in Theorem 1.3. The section is structured as follows: First, we derive an auxiliary estimate for the solution of a semidiscrete stochastic parabolic equation with non-homogeneous Dirichlet boundary conditions. This estimate plays a crucial role in adapting the Carleman estimate (1.6) to this setting. Second, we present a modified Carleman estimate adapted to the semidiscrete stochastic parabolic operator with nonhomogeneous boundary data, and provide its proof. Finally, we prove Theorem 1.3 by combining the Carleman estimate obtained in the second step with geometric arguments inspired by Yamamoto_2009 .
Lemma 3.1.
Let $\xi\in H_{\mathcal{F}}^{1}(0,T;H_{h}^{1/2}(\partial\mathcal{M})\cap H_{h}^{1}(\partial\mathcal{M}))$ with $\xi(0)=0$. Then, for a.e. $\omega\in\Omega$, the solution $u\in L^{2}(0,T;H_{h}^{2}(\mathcal{M}))\cap C(0,T;L_{h}^{2}(\mathcal{M}))$ of the system:
(3.1) $\begin{cases}du-\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}u)\,dt=0,&\text{in }\mathcal{M}\times(0,T),\\ u=\xi,&\text{on }\partial\mathcal{M}\times(0,T),\\ u(0)=0,&\text{in }\mathcal{M},\end{cases}$
satisfies the following stability estimate for any constant $C>0$:
(3.2) $\|u\|_{L^{2}_{\mathcal{F}}(0,T;H_{h}^{2}(\mathcal{M}))}+\|u\|_{L^{2}_{\mathcal{F}}(\Omega;C([0,T];L_{h}^{2}(\mathcal{M})))}\leq C\|\xi\|_{H^{1}_{\mathcal{F}}(0,T;H_{h}^{1/2}(\partial\mathcal{M})\cap H_{h}^{1}(\partial\mathcal{M}))}.$
Moreover, the following boundary estimate holds for $t_{r}^{i}(D_{i}u)$ on $\Gamma^{+}$:
(3.3) $\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}t_{r}^{i}(|D_{i}u|^{2})\,dt\leq C\|\xi\|_{H^{1}_{\mathcal{F}}(0,T;H_{h}^{1/2}(\partial\mathcal{M})\cap H_{h}^{1}(\partial\mathcal{M}))}.$
Proof.
Noting that for a.e. $\omega\in\Omega$ and $t\in[0,T]$, $\xi\in H_{h}^{1/2}(\partial\mathcal{M})$, there exists $v\in H_{h}^{1}(\mathcal{M})$ such that for a.e. $\omega\in\Omega$ and $t\in[0,T]$, $v=\xi$ on $\partial\mathcal{M}$ and
(3.4) $\|v\|_{H_{h}^{1}(\mathcal{M})}=\|\xi\|_{H_{h}^{1/2}(\partial\mathcal{M})}.$
Let $\tilde{u}=u-v$. Since $v(0)=0$ by (3.4) and $\xi(0)=0$, it follows that $\tilde{u}(0)=u(0)-v(0)=0$. Moreover, on $\partial\mathcal{M}\times(0,T)$ we have $\tilde{u}=\xi-\xi=0$. Therefore, $\tilde{u}$ satisfies
(3.5) $\begin{cases}d\tilde{u}-\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}\tilde{u})\,dt=\left(\partial_{t}v-\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}v)\right)\,dt,&\text{in }\mathcal{M}\times(0,T),\\ \tilde{u}=0,&\text{on }\partial\mathcal{M}\times(0,T),\\ \tilde{u}(0)=0,&\text{in }\mathcal{M}.\end{cases}$
Multiplying the first equation of (3.5) by $\tilde{u}$ and integrating over $\mathcal{M}$, we get
$\int_{\mathcal{M}}\tilde{u}\,d\tilde{u}-\int_{\mathcal{M}}\tilde{u}\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}\tilde{u})\,dt=\left(\int_{\mathcal{M}}\tilde{u}\,\partial_{t}v-\int_{\mathcal{M}}\tilde{u}\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}v)\right)\,dt.$
Applying integration by parts to the second integral on both sides, we obtain
$\int_{\mathcal{M}}\tilde{u}\,d\tilde{u}+\sum_{i=1}^{n}\int_{\mathcal{M}}\gamma_{i}|D_{i}\tilde{u}|^{2}\,dt=\left(\int_{\mathcal{M}}\tilde{u}\,\partial_{t}v+\sum_{i=1}^{n}\int_{\mathcal{M}}\gamma_{i}D_{i}v\,D_{i}\tilde{u}\right)\,dt,$
where we used $\tilde{u}=0$ on $\partial\mathcal{M}$. Observing that $\tilde{u}\,d\tilde{u}=\frac{1}{2}\,d(\tilde{u}^{2})$ and applying Young’s inequality to the right-hand side of the previous inequality, we get
$\frac{1}{2}\int_{\mathcal{M}}\,d(\tilde{u}^{2})+\frac{1}{2}\sum_{i=1}^{n}\int_{\mathcal{M}}\gamma_{i}|D_{i}\tilde{u}|^{2}\,dt\leq\frac{1}{2}\left(\int_{\mathcal{M}}|\tilde{u}|^{2}+\int_{\mathcal{M}}|\partial_{t}v|^{2}+\sum_{i=1}^{n}\int_{\mathcal{M}}\gamma_{i}|D_{i}v|^{2}\right)\,dt.$
Now, adding $\frac{1}{2}\|\tilde{u}\|_{L^{2}(\mathcal{M})}$ and using (3.4) in the above inequality, we obtain
(3.6) $\frac{1}{2}\,d\left(\|\tilde{u}\|^{2}_{L_{h}^{2}(\mathcal{M})}\right)+\frac{1}{2}\|\tilde{u}\|^{2}_{H_{h}^{1}(\mathcal{M})}\,dt\leq\left(\|\tilde{u}\|^{2}_{L_{h}^{2}(\mathcal{M})}+C\|\xi\|_{H^{1}(0,T;H_{h}^{1/2}(\partial\mathcal{M}))}\right)\,dt.$
Letting $E(t)=\|\tilde{u}\|^{2}_{L^{2}_{h}(\mathcal{M})}$ and $\mathcal{C}_{1}=C\|\xi\|_{H^{1}(0,T;H_{h}^{1/2}(\partial\mathcal{M}))}$ in (3.6), we have
$\frac{1}{2}\,dE(t)\leq(E(t)+\mathcal{C}_{1})\,dt.$
Multiplying by $e^{-2t}$ and using Itô’s lemma, we obtain
$\begin{split}e^{-2t}\,dE(t)-e^{-2t}E(t)\,dt\leq&e^{-2t}\mathcal{C}_{1}\,dt,\\ d\left(e^{-2t}E(t)\right)\leq&e^{-2t}\mathcal{C}_{1}\,dt.\end{split}$
Since $E(0)=0$, integrating over $(0,t)$ for $t\in[0,T]$, we have
(3.7) $\|\tilde{u}\|^{2}_{L_{h}^{2}(\mathcal{M})}=E(t)\leq\int_{0}^{t}e^{2(t-s)}\mathcal{C}_{1}\,ds=\frac{-1+e^{2t}}{2}\mathcal{C}_{1},$
for all $t\in[0,T]$. Recalling that $u=\tilde{u}+v$, and using (3.4) and (3.7), we get
$\|u\|^{2}_{L^{2}_{h}(\mathcal{M})}\leq\|\tilde{u}\|^{2}_{L_{h}^{2}(\mathcal{M})}+\|v\|^{2}_{L_{h}^{2}(\mathcal{M})}\leq\frac{e^{2t}-1}{2}\mathcal{C}_{1}+C\|\xi\|_{H^{1}(0,T;H_{h}^{1/2}(\partial\mathcal{M}))}.$
Therefore, taking the supremum over $[0,T]$ in the above inequality, we conclude that
$\|u\|_{C(0,T;L_{h}^{2}(\mathcal{M}))}\leq C\|\xi\|_{H^{1}(0,T;H_{h}^{1/2}(\partial\mathcal{M}))}.$
Furthermore, using (3.6), (3.7) and integrating over $(0,T)$, we have
$\|u\|_{L_{h}^{2}(0,T;H^{1}(\mathcal{M}))}\leq C\|\xi\|_{H_{h}^{1}(0,T;H^{1/2}(\partial\mathcal{M}))}.$
Combining the last two inequalities and taking the expectation, we derive the first estimate (3.2).
In turn, we note that multiplying by $x_{j}A_{j}D_{j}u$, where $x_{j}:=e_{j}\cdot x$, the main equation from (3.1) we obtain
(3.8) $du\,x_{j}A_{j}D_{j}u=\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}u)x_{j}D_{j}A_{j}udt.$
First, let us focus on the right-hand side above. Thanks to (1.13) we write
(3.9) $\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}u)x_{j}A_{j}D_{j}udt=$ $x_{j}D_{j}\gamma_{j}|A_{j}D_{j}u|^{2}+x_{j}A_{j}\gamma_{j}D_{j}^{2}u\,D_{j}A_{j}u$
$+\sum_{i\neq j}x_{j}D_{i}(\gamma_{i}D_{i}u)\,D_{j}A_{j}u.$
Then, writing $D_{j}^{2}uD_{j}A_{j}u=\frac{1}{2}D_{j}(|D_{j}u|^{2})$, integrating the results over $G\times(0,T)$, and applying an integration by parts it follows
$\frac{1}{2}\int_{\mathcal{M}}x_{j}A_{j}\gamma_{j}D_{j}(|D_{j}u|^{2})=$ $-\frac{1}{2}\int_{\mathcal{M}^{\ast}_{j}}D_{j}(A_{j}\gamma_{j}x_{j})|D_{j}u|^{2}$
$+\frac{1}{2}\int_{\partial_{j}\mathcal{M}}x_{j}A_{j}\gamma_{j}t_{r}^{j}(|D_{j}u|^{2})\nu_{j}.$
Thus, denoting $\Gamma_{j}:=\{x\in\partial_{j}\mathcal{M}\mid x_{j}\nu_{j}=1\}\subset\mathcal{M}_{j}$ it follows that the above equals
(3.10) $\frac{1}{2}\int_{\mathcal{M}}x_{j}A_{j}\gamma_{j}D_{j}(|D_{j}u|^{2})=-\frac{1}{2}\int_{\mathcal{M}^{\ast}_{j}}D_{j}(A_{j}\gamma_{j}x_{j})|D_{j}u|^{2}+\frac{1}{2}\int_{\Gamma_{j}}A_{j}\gamma_{j}t_{r}^{j}(|D_{j}u|^{2}).$
Hence, integrating over $\mathcal{M}\times(0,T)$ in (3.9), and then using equation (3.10) we obtain
(3.11) $\int_{G}\,x_{j}A_{j}D_{j}u\,du=$ $\int_{G}x_{j}D_{j}\gamma_{j}\,|A_{j}D_{j}u|^{2}-\frac{1}{2}\int_{\mathcal{M}_{j}^{\ast}\times(0,T)}D_{j}(A_{j}(\gamma_{j})x_{j})|D_{j}u|^{2}$
$+\frac{1}{2}\int_{\Gamma_{j}\times(0,T)}A_{j}\gamma_{j}t_{r}^{j}(|D_{j}u|^{2})+\sum_{i\neq j}\int_{G}x_{j}D_{i}(\gamma_{i}\,D_{i}u)\,D_{j}A_{j}u.$
Finally, using (3.8) in the integral of the left-hand side above, that $\gamma_{i}$ is bounded for each $j\in\llbracket 1,n\rrbracket$ and the Young inequality we conclude that there exists a constant $C$ independent of $h$ such that
(3.12) $\int_{\Gamma_{j}\times(0,T)}t_{r}^{j}(|D_{j}u|^{2})dt\leq C\left\|u\right\|^{2}_{L^{2}(0,T;H_{h}^{2}(\mathcal{M}))}.$
The inequality (3.12) stands for a bound for the elements of the boundary $\Gamma_{j}$, in the direction $x_{j}$, such that its normal function is equal to one. Similarly, it is possible to obtain a bound for the normal derivative on the boundary nodes where the normal function defined in (1.2) is minus one, that is, to adopt the previous methodology by using the term $(1-x_{j})A_{j}D_{j}$ instead of the used to obtain the equation (3.8). Moreover, we emphasize that we can obtain the inequality (3.12) in any direction. In conclusion, the solution of the system (3.1) satisfies
(3.13) $\sum_{j=1}^{n}\int_{\Gamma_{j}\times(0,T)}t_{r}^{j}(|D_{j}u|^{2})dt\leq C\left\|u\right\|^{2}_{L^{2}(0,T;H_{h}^{2}(\mathcal{M}))}.$
Overall, combining (3.2) and (3.13) we obtain (3.3). ∎
Remark 3.2.
We emphasize that the key argument of the proof of Lemma 3.1 is that we are dealing with square domains.
3.1. Semidiscrete Carleman estimate with nonhomogeneous Dirichlet condition.
In the Carleman estimate (1.10), the function $\psi$ is chosen as $|x-x^{\ast}|^{2}$. More generally, we may consider the following
(3.14) $\psi(x)=\Psi(x),$
where $|\nabla\Psi|>0$ in $G$, so that the assumptions required to obtain the Carleman estimate (1.10) remain valid. This generalized setting will later allow us to derive a Carleman estimate for a stochastic parabolic operator with non-homogeneous Dirichlet boundary conditions.
Based on this choice of $\psi$, we now state the corresponding Carleman estimate, which extends (1.10) to the present setting. For completeness, we also provide its proof in the following.
Theorem 3.3.
[…]
(3.15) $\begin{split}&J(w)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|w|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}w|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|g|^{2}dt\\ &\leq\,C_{s_{0},\varepsilon}\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}|f|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|D_{i}g|^{2}dt+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}st_{r}^{i}(e^{2s\varphi}|D_{i}w|^{2})\,dt\right.\\ &\left.+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|w|^{2}\right|_{t=T}\right)+\mathcal{C}\tau^{3}e^{\tau\mathcal{C}}\|\xi\|_{H^{1}_{\mathcal{F}}(0,T;H^{1}(\partial\mathcal{M})}.\end{split}$
for all $\tau\geq\tau_{0}$, $0<h\leq h_{0}$, $sh\leq\epsilon_{0}$, where $J(w)\triangleq\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi e^{2s\theta\varphi}\,|D_{i}w|^{2}\,dt+\mathbb{E}\int_{Q}s\varphi e^{2s\theta\varphi}|A_{i}D_{i}w|^{2}\,dt$, $f\in L^{2}_{\mathcal{F}}(0,T;L_{h}^{2}(\mathcal{M}))$, $g\in L^{2}_{\mathcal{F}}(0,T;H^{1}(\mathcal{M}))$,
$\xi\in H^{1}_{\mathcal{F}}(0,T;H_{h}^{1}(\partial\mathcal{M})\cap H_{h}^{1/2}(\partial\mathcal{M)})$ satisfying compatibility condition $\xi(0)=0$, and $w$ satisfies $dw-\sum_{i\in\llbracket 1,n\rrbracket}D_{i}(\gamma_{i}D_{i}w)\,dt=fdt+gdB(t)$ with $w=\xi$ in $\partial\mathcal{M}$ and $w(0)=0$ in $\mathcal{M}$.
Proof.
Let $u$ such that satisfies (3.1) and taking z=wuz=w-u, we obtain
[…]
Applying (1.10) to $z$, we have for all $\tau\geq\tau_{0}$, $0<h\leq h_{0}$, $sh\leq\epsilon_{0}$,
$\begin{split}&J(z)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|z|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}z|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|g|^{2}dt\\ &\leq\,C_{s_{0},\varepsilon}\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}|f|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|D_{i}g|^{2}dt+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}st_{r}^{i}(e^{2s\varphi}|D_{i}z|^{2})\,dt\right.\\ &\left.+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|z|^{2}\right|_{t=T}\right),\end{split}$
where $J(z)\triangleq\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi e^{2s\theta\varphi}\,|D_{i}z|^{2}\,dt+\mathbb{E}\int_{Q}s\varphi e^{2s\theta\varphi}|A_{i}D_{i}z|^{2}\,dt$. Recalling that $w=z+u$ and using the triangular inequality, we can obtain that
$\begin{split}&J(w)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|w|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}w|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|g|^{2}dt\\ &\leq\,C_{s_{0},\varepsilon}\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}|f|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|D_{i}g|^{2}dt+\sum_{i=1}^{n}\mathbb{E}\int_{\Gamma^{+}}st_{r}^{i}(e^{2s\varphi}|D_{i}w|^{2})\,dt\right.\\ &\left.+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|w|^{2}\right|_{t=T}+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}st_{r}^{i}(e^{2s\varphi}|D_{i}u|^{2})\,dt+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|u|^{2}\right|_{t=T}\right)\\ &+J(u)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|u|^{2}\,dt+\sum_{i,j\in\llbracket 1,n\rrbracket}\mathbb{E}\int_{Q}s^{-1}e^{2s\varphi}|D_{ij}^{2}u|^{2}\,dt.\end{split}$
Applying el Lemma 3.1 and noting that
$\begin{split}J(u)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|u|^{2}&\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}u|^{2}\,dt+\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|u(T)|^{2}\\ &\leq\,\mathcal{C}\tau^{3}e^{\tau\mathcal{C}}\left(\|u\|_{L^{2}_{\mathcal{F}}(\Omega;L^{2}((0,T);H_{h}^{2}(\mathcal{M}))}+\|u(T)\|_{L^{2}_{\mathcal{F}_{T}}(\Omega;L_{h}^{2}(\mathcal{M})}\right),\end{split}$
we get the desired result. ∎
Based on the Carleman estimate with non-homogeneous boundary condition (3.15), we are now in a position to prove the main stability result of this section, namely Theorem 1.3.
3.2. Proof of Theorem 1.3.
In order to estimate the solution of (1.8) in $\mathcal{M}_{0}\times(\varepsilon,T-\varepsilon)$ by Cauchy data at $\Gamma^{+}$, we need to choose a suitable weight function $\varphi$. For this, we can choose a bounded domain $G_{1}$ with a smooth boundary such that
$G\subsetneq G_{1},\quad\overline{\Sigma}\triangleq\overline{\partial G\cap G_{1}},\quad\partial G\setminus\Sigma\subset\partial G_{1}.$
we can see that $G_{1}$ is the union of $G$ with other set $\tilde{G}$ such that $\partial\tilde{G}\cap\overline{G}=\Sigma$ and that $G_{1}\setminus\overline{G}$ have a non-empty open set.
$\begin{split}\Psi(x)>0,x\in G_{1}\quad&\Psi(x)=0,\,x\in\partial G_{1},\\ |\nabla\Psi(x)|>0,\;&x\in\overline{G_{1}}\cap\overline{G}.\end{split}$
Then, since $\overline{G}_{0}\subset G_{1}$, we can choose a sufficiently large $N>1$ such that
(3.16) $\left\{x\in G_{1}:\Psi(x)>\frac{4}{N}\|\Psi\|_{C(\overline{G_{1}})}\right\}\cap\overline{G}\supset G_{0}.$
Moreover, we choose $\beta>0$ such that
(3.17) $2\beta\varepsilon^{2}>\|\Psi\|_{C(\overline{G_{1}})}>\beta\varepsilon^{2}$
We arbitrarily fix $t_{0}\in[\sqrt{2}\varepsilon,T-\sqrt{2}\varepsilon]$. We set $\varphi(x)s(t)=\tau e^{\lambda\psi(x)-\lambda\theta(t)}$ with fixed large parameter $\lambda>0$ and $\psi(x)-\theta(t)=\Psi(x)-\beta(t-t_{0})^{2}$, $\mu_{k}=\exp{\left(\lambda\left(\frac{k}{N}\|\Psi\|_{C(\overline{G_{1}})}-\frac{\beta\varepsilon^{2}}{N}\right)\right)}$, $k=1,2,3,4$, and
(3.18) $G^{(k)}\triangleq\left\{(x,t):x\in\overline{G},\Phi(x,t)>\mu_{k}\;\text{such that}\;\Phi\triangleq e^{\lambda\psi(x)-\lambda\theta(t)}\right\}.$
In the same way as in the proof of Theorem 5.1 in Yamamoto_2009 it can be verified that
(3.19) $G_{0}\times\left(t_{0}-\frac{\varepsilon}{\sqrt{N}},t_{0}+\frac{\varepsilon}{\sqrt{N}}\right)\subset G^{(k)}\subset\overline{G}\times(t_{0}-\sqrt{2}\varepsilon,t_{0}+\sqrt{2}\varepsilon)$
and
(3.20) $\begin{split}\partial G^{(1)}\subset\Sigma^{(1)}\cup\Sigma^{(2)},&\\ \Sigma^{(1)}\subset\Sigma\times(0,T),\quad&\quad\Sigma^{(2)}\triangleq\{(x,t):x\in G,\Phi(x,t)=\mu_{1}\}.\end{split}$
Let $\mathcal{M}^{(k)}\triangleq G^{(k)}\cap\mathcal{M}$. In order to apply Theorem 3.3, we need a cut-off function because we have no data on $\partial G^{(1)}\setminus(\Sigma\times(0,T))$. Let $\chi\in C^{\infty}(\overline{G})$ such that $0\leq\chi\leq 1$ and
(3.21) $\chi(x,t)\triangleq\left\{\begin{array}[]{cc}1&\Phi(x,t)>\mu_{3},\\ 0&\Phi(x,t)<\mu_{2}.\end{array}\right$.
We set $v=\chi w$, and we have
$dv-\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}v)dt=\left(\sum_{i=1}^{n}a_{1i}A_{i}D_{i}(v)+a_{2}v+\mathbf{F}\right)dt+a_{3}vdB(t),\quad\text{in}\,Q$
where
$\begin{split}\mathbf{F}\triangleq&\chi_{t}w-\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}(\chi)A(w))+A_{i}D_{i}(\chi)A_{i}(\gamma_{i}D_{i}w)+a_{1i}A_{i}D_{i}(\chi)A_{i}^{2}(w)\\ &+\frac{h^{2}}{4}\left(D_{i}^{2}(\chi)D_{i}(\gamma_{i}D_{i}(w))+a_{1i}D_{i}^{2}(\chi)A_{i}D_{i}(w)+2a_{1i}A_{i}D_{i}(\chi)D_{i}^{2}(w)\right).\end{split}$
By (3.20) and (3.21), we see that for all $i\in\llbracket 1,n\rrbracket$
(3.22) $v=D_{i}v=0\quad\quad\text{on}\,\,\Sigma^{(2)}\cap\partial\mathcal{M}.$
Since $t_{0}\in[\sqrt{2}\varepsilon,T-\sqrt{2}\varepsilon]$, we obtain the following result
$\max\left\{\Phi(x,0),\Phi(x,T)\right\}\leq\Psi(x)-2\beta\varepsilon^{2}\leq 0,$
which implies $(\mathcal{M}\times\{t=0\})\cup(\mathcal{M}\times\{t=T\})\subset\sup{(\chi)}$ and then
(3.23) $v(0)=v(T)=0.$
Applying (3.15) on $v$ and taking into account (3.22), (3.23), the choice of $\Sigma$ is arbitrary, so it is made to coincide $\Sigma\cap\partial\mathcal{M}$ with $\Gamma^{+}$ and taking $\eta_{i}$ as $t_{r}^{i}(D_{i}w)$ on $\Gamma^{+}$, we have for all $\tau\geq\tau_{0}$, $0<h\leq h_{0}$, $sh\leq\epsilon_{0}$
$\begin{split}&J(v)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|v|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}v|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|a_{3}v|^{2}dt\\ &\leq\,C_{s_{0},\varepsilon}\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}\left|\sum_{i=1}^{n}a_{1i}A_{i}D_{i}(v)+a_{2}v+\mathbf{F}\right|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|D_{i}(a_{3}v)|^{2}dt\right.\\ &\left.+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}st_{r}^{i}(e^{2s\varphi})|\eta_{i}|^{2}\,dt\right)+\mathcal{C}\tau^{3}e^{\tau\mathcal{C}}\|\xi\|^{2}_{H^{1}_{\mathcal{F}}(0,T;H_{h}^{1}(\Gamma^{+})},\end{split}$
where $J(v)\triangleq\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi e^{2s\theta\varphi}\,|D_{i}v|^{2}\,dt+\mathbb{E}\int_{Q}s\varphi e^{2s\theta\varphi}|A_{i}D_{i}v|^{2}\,dt$.
Using the Young inequality and (1.13) to the first and second integral from the right-hand side above, respectively, we can bound the above inequality as
(3.24) $\begin{split}J(v)+&\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|v|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}v|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|a_{3}v|^{2}dt\\ &\leq\,\mathcal{C}\tau^{3}e^{\tau\mathcal{C}}\|\xi\|^{2}_{H^{1}_{\mathcal{F}}(0,T;H_{h}^{1}(\Gamma^{+})}+C\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}|\mathbf{F}|^{2}dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q}\,e^{2s\varphi}|A_{i}D_{i}v|^{2}\,dt\right.\\ &\left.+\mathbb{E}\int_{Q}e^{2s\varphi}|a_{2}v|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}\left(|A_{i}v|^{2}+|D_{i}v|^{2}\right)\,dt\right.\\ &\left.+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}st_{r}^{i}(e^{2s\varphi})|\eta_{i}|^{2}\,dt\right).\end{split}$
Noting that using (2.5), (1.12) and $e^{-2s\varphi}A_{i}(e^{2s\varphi})=1+\mathcal{O}((sh)^{2})$, for each $i\in\llbracket 1,n\rrbracket$ we have
$\begin{split}\mathbb{E}\int_{Q_{i}^{\ast}}e^{2s\varphi}|A_{i}v|^{2}\,dt\leq&\mathbb{E}\int_{Q_{i}^{\ast}}e^{2s\varphi}A_{i}(|v|^{2})\,dt=\;\mathbb{E}\int_{Q}A_{i}(e^{2s\varphi})|v|^{2}\,dt+\frac{h}{2}\mathbb{E}\int_{\partial_{i}Q}|v|^{2}t_{r}^{i}(e^{2s\varphi})\,dt\\ \leq&\mathbb{E}\int_{Q}e^{2s\varphi}|v|^{2}\,dt+\frac{h}{2}\mathbb{E}\int_{\partial_{i}Q}|v|^{2}t_{r}^{i}(e^{2s\varphi})\,dt+\mathcal{O}((sh)^{2})\int_{Q}e^{2s\varphi}|v|^{2}\,dt\end{split}$
Therefore, combining the above inequalities and (3.24), we obtain the following
$\begin{split}&J(v)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|v|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}v|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|a_{3}v|^{2}dt\\ &\leq\,C\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}|\mathbf{F}|^{2}dt+\mathbb{E}\int_{Q}\,e^{2s\varphi}|v|^{2}\,dt+\mathcal{O}((sh)^{2})\int_{Q}e^{2s\varphi}|v|^{2}\,dt\right.\\ &+\sum_{i=1}^{n}\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}|A_{i}D_{i}v|^{2}+\frac{h}{2}\int_{0}^{T}\int_{\Gamma^{+}}|v|^{2}t_{r}^{i}(e^{2s\varphi})+\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|D_{i}v|^{2}\right)\,dt\\ &\left.+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}st_{r}^{i}(e^{2s\varphi})|\eta_{i}|^{2}\,dt+\mathbb{E}\int_{Q}e^{2s\varphi}|v|^{2}\,dt\right)+\mathcal{C}\tau^{3}e^{\tau\mathcal{C}}\|\xi\|^{2}_{H^{1}_{\mathcal{F}}(0,T;H_{h}^{1}(\Gamma^{+}))}.\end{split}$
Recalling that $s(t)=\tau\theta(t)$ and letting $\tau$ sufficiently large and taking a convenient constant, we notice that the second, third, and fifth integrals are absorbed by the left side in the above inequality, and the fourth integral is controlled by the last integral on the right side. Thus, the above inequality can be rewritten as follows
[…]
Now, by (3.19), we have $\overline{G}_{0}\times\left(t_{0}-\frac{\varepsilon}{\sqrt{N}},t_{0}+\frac{\varepsilon}{\sqrt{N}}\right)\subset G^{(4)}$ and then
(3.26) $\overline{\mathcal{M}_{0}}\times\left(t_{0}-\frac{\varepsilon}{\sqrt{N}},t_{0}+\frac{\varepsilon}{\sqrt{N}}\right)\subset\mathcal{M}^{(4)}\subset\mathcal{M}$
for sufficiently small $h\in(0,h_{0})$. Then, observing (3.21) and (3.26) we see that
(3.27) $\begin{split}&J(v)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|v|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}v|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|a_{3}v|^{2}dt\\ &\geq\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|v|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi e^{2s\theta\varphi}\,|D_{i}v|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D^{2}_{ij}v|^{2}\,dt\\ &\geq C\tau^{-1}e^{2\mu_{4}\tau}\int_{t_{0}-\frac{\varepsilon}{\sqrt{N}}}^{t_{0}+\frac{\varepsilon}{\sqrt{N}}}\left(\mathbb{E}\int_{\mathcal{M}_{0}}|w|^{2}+\sum_{i=1}^{n}\mathbb{E}\int_{(\mathcal{M}_{0})_{i}^{\ast}}\,|D_{i}w|^{2}+\sum_{i,j=1}^{n}\mathbb{E}\int_{(\mathcal{M}_{0})_{ij}^{\ast}}|D^{2}_{i}w|^{2}\right)\,dt.\end{split}$
On the other hand, we can verify that there exists a sufficiently small $h_{1}$ such that
(3.28) $\partial_{t}\chi=A_{i}D_{i}\chi=D_{i}^{2}\chi=0\quad(x,t)\in Q^{{}^{\prime}}$
for all $h\in(0,h_{1})$, where
$Q^{{}^{\prime}}\triangleq\left\{(x,t)\in Q\;:\quad\Phi(x,t)>\mu_{3}+\frac{1}{2}(\mu_{4}-\mu_{3})\right\}.$
We apply Proposition 1.8 and (3.28) to yield
(3.29) $\begin{split}&\mathbb{E}\int_{Q}e^{2s\varphi}|\mathbf{F}|^{2}\,dt\\ &\leq Ce^{2\tau\left(\mu_{3}+\frac{1}{2}(\mu_{4}-\mu_{3})\right)}\mathbb{E}\int_{Q\setminus Q^{{}^{\prime}}}\left(|w|^{2}+\sum_{i=1}^{n}\left[\left(1+\frac{h^{4}}{16}\right)|A_{i}D_{i}w|^{2}+\frac{h^{4}}{16}|D_{i}^{2}w|^{2}\right]\right)\,dt\\ &\leq Ce^{\tau(\mu_{4}+\mu_{3})}\|w\|^{2}_{L_{\mathcal{F}}^{2}(0,T;H_{h}^{2}(\mathcal{M}))},\end{split}$
where in the second term of the second line, we have used that
$\mathbb{E}\int_{Q}|A_{i}D_{i}w|^{2}\,dt\leq\mathbb{E}\int_{Q}A_{i}(|D_{i}w|^{2})\,dt\leq\mathbb{E}\int_{Q_{i}^{\ast}}|D_{i}w|^{2}\,dt,$
which follows after the combination of (2.5) and (1.12). Therefore, from (3.25), (3.27) and (3.29) it follows that
[…]
which implies
[…]
where we have used that there exists a sufficiently large $\tau_{1}$ such that
$C\tau e^{-\tau(\mu_{4}-\mu_{3})}\leq e^{-\frac{1}{2}\tau(\mu_{4}-\mu_{3})}$
for all $\tau_{1}<\tau<C\varepsilon_{0}h^{-1}$.
Consider $f(\tau)=ae^{-\frac{1}{2}\tau(\mu_{4}-\mu_{3})}+be^{2\mathcal{C}\tau}$ with $\tau_{1}<\tau<C\varepsilon_{0}h^{-1}$, $a=\|w\|^{2}_{L^{2}_{\mathcal{F}}{(0,T;H_{h}^{2}(\mathcal{M}))}}$ and $b=\|\eta\|^{2}_{L^{2}_{\mathcal{F}}(0,T;L_{h}^{2}(\Gamma^{+}))}+\|\xi\|^{2}_{H^{1}_{\mathcal{F}}(0,T;H_{h}^{1}(\Gamma^{+}))}.$ As $\mu_{4}-\mu_{3}>0$, then from Lemma 3.39 in LOPD the function $f$ satisfies the following.
$\min_{\tau\in[\tau_{1},C\varepsilon_{0}h^{-1}]}f(\tau)\leq 2\max\left\{be^{2\mathcal{C}\tau_{1}},ae^{-\frac{1}{2h}C\varepsilon_{0}(\mu_{4}-\mu_{3})},ae^{-\frac{1}{2}\tau_{\ast}(\mu_{4}-\mu_{3})}\right\},$
for any $\tau_{\ast}\leq\frac{2}{4\mathcal{C}+(\mu_{4}-\mu_{3})}\ln{\left(\frac{a}{b}\right)}$. Given that (3.30) is valid for all $\tau\in[\tau_{1},C\varepsilon_{0}h^{-1}]$ and taking $\tau_{\ast}=\frac{2}{4\mathcal{C}+(\mu_{4}-\mu_{3})}\ln{\left(\frac{a}{b}\right)}$, it follows in particular that:
(3.31) $\begin{split}\|w\|^{2}_{L^{2}_{\mathcal{F}}\left((t_{0}-\frac{\varepsilon}{\sqrt{N}},t_{0}+\frac{\varepsilon}{\sqrt{N}};H_{h}^{2}(\mathcal{M}_{0})\right)}\leq\;2\max\left\{be^{2\mathcal{C}\tau_{1}},ae^{-\frac{1}{2h}\mathcal{C}\varepsilon_{0}(\mu_{4}-\mu_{3})},a^{\kappa}b^{1-\kappa}\right\}\end{split}$
with
$\kappa\triangleq\frac{4\mathcal{C}}{4\mathcal{C}+(\mu_{4}-\mu_{3})}\in(0,1)$
Finally, for the estimate in (3.31), we choose points $t_{0}=\sqrt{2}\varepsilon+\frac{j\varepsilon}{\sqrt{N}}$, for $j=0,1,2,\ldots,m$ where $m$ is chosen such that
$\sqrt{2}\varepsilon+\frac{m\varepsilon}{\sqrt{N}}\leq T-\sqrt{2}\leq\sqrt{2}\varepsilon+\frac{m\varepsilon}{\sqrt{N}}$
Summing the estimates over $j$ from 0 to $m$ and then replacing $\varepsilon$ by $\sqrt{2}\varepsilon$ (which is permissible since $\varepsilon>0$ is arbitrary) produces the desired inequality (1.9), thus completing the proof of Theorem 1.3.
Acknowledgments
R. Lecaros was partially supported by FONDECYT (Chile) Grant 1221892 and Proyecto Interno USM 2025 PI_\_LIR_\_25_\_14. A. A. Pérez acknowledges the support of Vicerrectoría de Investigación y postgrado, Universidad del Bío-Bío, proyect IN2450902 and FONDECYT Grant 11250805. M. F. Prado gratefully acknowledges the support of the Institutional Scholarship Fund of the Universidad de Valparaiso (FIB-UV) and Proyecto Interno USM 2025 PI_\_LIR_\_25_\_14.
Appendix A Calculus results related to the weight functions and Discrete spaces
This section is devoted to presenting key results on discrete operations and discrete spaces, which are essential for the proof of the Carleman estimate. Estimates for the Carleman weight functions in arbitrary dimensions have been established in LOP-2020 . For convenience, we set $r=e^{s\varphi}$ and $\rho=r^{-1}$, where $\varphi(t,x)=\theta(t)\psi(x)$, and consider $s$ and $h$ to be positive parameters with $s$ large and $h$ small. We are particularly interested in the dependence on $s$ and $h$ in the following basic estimates over the domain $G$.
Concerning discrete spaces, we also recall a Sobolev-type inequality from EstimateCarlemanStochasticControl , which will play a key role in controlling norms when switching between discrete spaces.
We assume $s\geq 1$ throughout this section. We begin by presenting two fundamental estimates for the weight function.
Lemma A.1 ((boyer-2010-elliptic, , Lemma 3.7)).
Let $\alpha$ and $\beta$ multi-indices. We have
(A.1) $\begin{split}\partial^{\beta}(r\partial^{\alpha}\rho)=&|\alpha|^{|\beta|}(-s\varphi)^{|\alpha|}\lambda^{|\alpha+\beta|}(\nabla\psi)^{\alpha+\beta}+|\alpha||\beta|(s\varphi)^{|\alpha|}\lambda^{|\alpha+\beta|-1}\mathcal{O}(1)+s^{|\alpha|-1}|\alpha|(|\alpha|-1)\mathcal{O}_{\lambda}(1)\\ =&\mathcal{O}_{\lambda}(s^{|\alpha|}).\end{split}$
Let $\sigma\in[-1,1]$ and $i\in\llbracket 1,d\rrbracket$. We have
(A.2) $\begin{split}\partial^{\beta}\left(r(x)(\partial^{\alpha}\rho)(x+\sigma h_{i})\right)=\mathcal{O}_{\lambda,\varepsilon}(s^{|\alpha|}(1+(sh)^{|\beta|}))e^{\mathcal{O}_{\lambda}(sh)}.\end{split}$
Provided $sh\leq\varepsilon$ we have $\partial^{\beta}\left(r(x)(\partial^{\alpha}\rho)(x+\sigma h_{i})\right)=\mathcal{O}_{\lambda,\varepsilon}(s^{|\alpha|})$. The same expressions hold with $r$ and $\rho$ interchanged and with $s$ changed to s-s.
Another useful estimate is the following
Corollary A.2 ((boyer-2010-elliptic, , Corollary 3.8)).
Let $\alpha,\beta$ and $\delta$ multi-indices. We have
(A.3) $\begin{split}\partial^{\delta}\left(r^{2}\,\partial^{\alpha}\rho\,\partial^{\beta}\rho\right)=&|\alpha+\beta|^{|\delta|}(-s\varphi)^{|\alpha+\beta|}\lambda^{|\alpha+\beta+\delta|}(\nabla\psi)^{\alpha+\beta+\delta}\\ &+|\delta||\alpha+\beta|(s\varphi)^{|\alpha+\beta|}\lambda^{|\alpha+\beta+\delta|-1}\mathcal{O}(1)\\ &+s^{|\alpha+\beta|-1}\left(|\alpha|(|\alpha|-1)+|\beta|(|\beta|-1)\right)\mathcal{O}(1)\\ &=\mathcal{O}_{\lambda}(s^{|\alpha+\beta|}).\end{split}$
Mimicking the notation for multi-index in the continuous setting, we define for the 2-dimensional index $k=(k_{i},k_{j})$ with $k_{i},k_{j}\in\mathbb{N}$ the discrete operator $\mathbf{D}_{h}^{k}:=D_{i}^{k_{i}}D_{j}^{k_{j}}$. Analogously, for the average operator, we define $\mathbf{A}_{h}^{l}:=A_{i}^{l_{i}}A_{j}^{l_{j}}$ with $l:=(l_{i},l_{j})$ being a two-dimensional index.
Proposition A.3 ((AA:perez:2024, , Proposition 3.3)).
Let $\alpha$ be a multi-index and $k,l$ be two-dimensional multi-indices, respectively. Let $i,j\in\llbracket 1,d\rrbracket$; provided that $sh\leq\varepsilon$, we have
(A.4) $r\mathbf{A}_{h}^{l}\mathbf{D}_{h}^{k}\partial^{\alpha}\rho$ $=r\partial^{k+\alpha}\rho+s^{|\alpha|+k_{i}}\mathcal{O}_{\lambda}((sh)^{2})$
$\quad+s^{|\alpha|+k_{j}}\mathcal{O}_{\lambda}((sh)^{2})+s^{|\alpha+k|}\mathcal{O}_{\lambda}((sh)^{2}).$
Another useful estimate is the following lemma
Lemma A.4 ((AA:perez:2024, , Lemma 3.4)).
Let $\alpha$, $\beta$ and $\delta$ be multi-indices and let $k,l$ be two-dimensional multi-indices. Provided $sh\leq 1$, we have
(A.5) $\mathbf{A}_{h}^{l}\mathbf{D}_{h}^{k}\left(\partial^{\beta}\left(r\partial^{\alpha}\rho\right)\right)=\partial_{i}^{k_{i}}\partial_{j}^{k_{j}}\partial^{\beta}\left(r\partial^{\alpha}\rho\right)+h^{2}\mathcal{O}_{\lambda}(s^{|\alpha|}).$
Moreover, for $i\neq j\in\llbracket 1,d\rrbracket$ we have
(A.6) $\mathbf{A}_{h}^{l}\mathbf{D}_{h}^{k}\partial^{\delta}\left(r^{2}\left(\partial^{\alpha}\rho\right)\partial^{\beta}\rho\right)=\partial^{k+\delta}\left(r^{2}(\partial^{\alpha}\rho)\partial^{\beta}\rho\right)+h^{2}\mathcal{O}(s^{|\alpha|+|\beta|}),$
and for $\sigma,\sigma^{\prime}\in[0,1]$ we get
(A.7) $\mathbf{A}_{h}^{l}\mathbf{D}^{k}_{h}\partial^{\beta}\left(r(x)\partial^{\alpha}\rho(x+\sigma he_{i}\right)$ $=\mathcal{O}_{\lambda}(s^{|\alpha|}),$
(A.8) $\mathbf{A}_{h}^{l}\mathbf{D}_{h}^{k}\partial^{\delta}\left(r^{2}(x)\left(\partial^{\alpha}\rho(x+\sigma he_{i})\right)\partial^{\beta}\rho(x+\sigma^{\prime}he_{j})\right)$ $=\mathcal{O}_{\lambda}(s^{|\alpha|+|\beta|}).$
Moreover, we present estimates for various compositions of the discrete average and derivative operators applied to the Carleman weight function. These results will play a fundamental role in the proof of the semidiscrete Carleman estimate.
Theorem A.5 ((AA:perez:2024, , Theorem 3.5)).
Let $\alpha$ be a multi-index and $l,k,n,m,p$ and $q$ two-dimensional multi-indices. For $i,j\in\llbracket 1,d\rrbracket$, and for $sh\leq 1$, we have
(A.9) $\mathbf{A}_{h}^{l}\mathbf{D}_{h}^{k}\partial^{\alpha}(r\mathbf{A}_{h}^{m}\mathbf{D}_{h}^{n}\rho)=$ $\partial^{n+\alpha}(r\partial^{n}\rho)++s^{|\alpha|+n_{i}}\mathcal{O}_{\lambda}((sh)^{2})$
$+s^{|\alpha|+n_{j}}\mathcal{O}_{\lambda}((sh)^{2})+s^{|n|}\mathcal{O}_{\lambda}((sh)^{2})$
and
(A.10) $\mathbf{A}_{h}^{l}\mathbf{D}_{h}^{k}\partial^{\beta}(r^{2}\mathbf{A}^{p}_{h}\mathbf{D}^{q}_{h}(\partial^{\alpha}\rho)\mathbf{A}_{h}^{m}\mathbf{D}_{h}^{n}\rho)=$ $\partial^{k+\beta}(r^{2}(\partial^{q+\alpha}\rho)\partial^{n}\rho)$
$+s^{|\alpha+n+q|}\mathcal{O}_{\lambda}((sh)^{2}).$
Finally, we conclude this section with a discrete Sobolev-type inequality, which plays a key role in controlling norms when transferring between discrete spaces.
Theorem A.6 ((EstimateCarlemanStochasticControl, , Theorem 1.5)).
Let $n>1$. There exists a constant $C$ independent of $h$ such that the following inequality holds:
$\|u\|_{L_{h}^{p^{\ast}}(\mathcal{M)}}\leq C\|u\|_{W_{h}^{1,p}(\mathcal{M})}\quad\forall u\in W_{h}^{1,p}(\mathcal{M}),$
with $u=0$ on $\partial\mathcal{M}$, where
  • If $1\leq p<n$, then $p^{\ast}$ is given by $\frac{1}{p^{\ast}}=\frac{1}{p}-\frac{1}{n}$.
  • If $p=n$ then $p^{\ast}\in[p,\infty)$.
Appendix B Proof of Semidiscrete Carleman estimate with homogeneous Dirichlet condition (Theorem 1.6).
In this section, we prove the semidiscrete Carleman estimate under homogeneous Dirichlet conditions. The argument follows a classical approach (see fursikov-1996 ), based on conjugating the original operator with a suitably chosen exponential weight. Our estimate is closely related to the one presented in boyer-2010-1d-elliptic , and we adopt its methodology whenever possible. The main difference is that we do not impose any boundary condition on the state or on the weight function, which leads to additional boundary terms that will be handled later on.
The proof is divided into three steps. First, in Section B.1, we decompose the conjugated operator into its symmetric part $P_{1}$, antisymmetric part $P_{2}$, and an auxiliary term $R$, which helps simplify the calculations. Next, in Section B.2, we estimate the cross-inner products between these operators. Finally, in Section B.3, we return to the original variable to conclude the argument.
B.1. Conjugated operator
For simplicity of notation, we write $r:=e^{s\varphi}$ and $\rho:=r^{-1}$, where our weight function is defined as $\varphi=e^{\lambda\psi}$, for $s\geq 1$, with $\psi\in C^{k}$ for $k$ sufficiently large and $\lambda\geq 1$. The proof of some technical results can be found in Section 5.
For all $i\in\llbracket 1,n\rrbracket$, let us consider the functions $\gamma_{i}$ such that $\mbox{reg}(\gamma)<\mbox{reg}^{0}$ and the following notation
$\nabla_{\gamma}f\triangleq(\sqrt{\gamma_{1}}D_{1}f_{1},\cdots,\sqrt{\gamma_{n}}D_{n}f_{n})\quad\text{and}\,\quad\Delta_{\gamma}f\triangleq\sum_{i=1}^{n}\gamma_{i}\partial_{x_{i}}^{2}f.$
Let $\mathcal{P}(w)\triangleq\,dw-\sum_{i=1}^{n}D_{i}\left(\gamma_{i}D_{i}w\right)dt=fdt+gdB(t)$. With respect to this weight function, the first step is to consider the change of variable $w=\rho z$. Our first task is to split the conjugate operator
$r\mathcal{P}(\rho z)=rd(\rho z)-r\sum_{i=1}^{n}D_{i}\left(\gamma_{i}D_{i}(\rho z)\right)dt$
into simple terms that we will estimate separately. Using Ito’s formula (cf. (klebaner2012introduction, , Theorem 4.18)) and notice that $r\rho=1$, we obtain the following result
$rd(\rho z)=dz+r\partial_{t}(\rho)z\,dt$
By (1.13) from Proposition 1.8, we have
$D_{i}\left(\gamma_{i}D_{i}(\rho z)\right)=$ $D_{i}(\gamma_{i}D_{i}\rho)A_{i}^{2}z+A_{i}(\gamma_{i}D_{i}\rho)D_{i}A_{i}z+D_{i}A_{i}\rho A_{i}(\gamma_{i}D_{i}z)+A_{i}^{2}\rho D_{i}(\gamma_{i}D_{i}z).$
Then, applying Proposition 1.8 and using $A_{i}\gamma_{i}=\gamma_{i}+h\mathcal{O}(1)$, we have
$A_{i}(\gamma_{i}D_{i}z)$ $=A_{i}\gamma_{i}A_{i}D_{i}z+\frac{h^{2}}{4}D_{i}\gamma_{i}D_{i}^{2}z=\gamma_{i}\,A_{i}D_{i}z+\frac{h^{2}}{4}D_{i}\gamma_{i}\,D^{2}_{i}z+h\mathcal{O}(1)\,A_{i}D_{i}z,$
$A_{i}(\gamma_{i}D_{i}\rho)$ $=A_{i}\gamma_{i}A_{i}D_{i}\rho+\frac{h^{2}}{4}D_{i}\gamma_{i}D_{i}^{2}\rho=\gamma_{i}\,A_{i}D_{i}\rho+\frac{h^{2}}{4}D_{i}\gamma_{i}\,D_{i}^{2}\rho+h\mathcal{O}(1)\,A_{i}D_{i}\rho,$
$D_{i}(\gamma_{i}D_{i}\rho)$ $=D_{i}\gamma_{i}A_{i}D_{i}\rho+A_{i}\gamma_{i}D_{i}^{2}\rho=D_{i}\gamma_{i}\,A_{i}D_{i}\rho+\gamma_{i}\,D_{i}^{2}\rho+h\mathcal{O}(1)\,D_{i}^{2}\rho.$
Thus, the conjugate operator can be written as
(B.1) $\begin{split}rd(\rho z)-r\sum_{i=1}^{n}D_{i}\left(\gamma_{i}D_{i}(\rho z)\right)dt+R_{h}z\,dt=&B_{1}z+B_{2}z\,dt+C_{1}z\,dt+C_{2}z\,dt+C_{3}z\,dt,\end{split}$
where
$\begin{split}C_{1}z:=&-\sum_{i=1}^{n}rA_{i}^{2}\rho\,D_{i}(\gamma_{i}D_{i}z),\ C_{2}z:=-\sum_{i=1}^{n}\gamma_{i}rD_{i}^{2}\rho A_{i}^{2}z,\ C_{3}z:=r\partial_{t}(\rho)z\end{split}$
$\begin{split}B_{1}z:=&dz,\ B_{2}z=-2\sum_{i=1}^{n}\gamma_{i}rD_{i}A_{i}\rho D_{i}A_{i}z;\end{split}$
and
$\begin{split}R_{h}z\triangleq\,&\sum_{i=1}^{n}\left(h\mathcal{O}(1)rD_{i}^{2}\rho+D_{i}\gamma_{i}rA_{i}D_{i}\rho\right)A_{i}^{2}z+\sum_{i=1}^{n}h_{i}\mathcal{O}(1)rD_{i}A_{i}\rho D_{i}A_{i}z\\ &+\sum_{i=1}^{n}\frac{h^{2}}{4}D_{i}\gamma_{i}rD_{i}^{2}\rho D_{i}A_{i}z+\sum_{i=1}^{n}\frac{h^{2}}{4}D_{i}\gamma_{i}\,rD_{i}A_{i}\rho D_{i}^{2}z.\end{split}$
Therefore, adding the terms
$C_{4}z\triangleq\,-\frac{h^{2}}{4}\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}(rD_{i}^{2}\rho)A_{i}z),\quad C_{5}z\triangleq\,-\frac{h^{2}}{4}\sum_{i=1}^{n}D_{i}(D_{i}(\gamma_{i}rD^{2}_{i}\rho)A_{i}z)$
$C_{6}z=-B_{3}z\triangleq\,-2s\Delta_{\gamma}\varphi\,z$
in (B.1) we have the following identity in $Q$
(B.2) $r\mathcal{P}(\rho z)=Cz\,dt+Bz-M_{h}z\,dt,$
where $Cz=C_{1}z+C_{2}z+C_{3}z+C_{4}z+C_{5}z+C_{6}z$, $Bz:=B_{1}z+(B_{2}z+B_{3}z)\,dt$ and $M_{h}z=C_{4}z+C_{5}z+R_{h}z$.
Thus, our next task is to estimate the cross-product $2\mathbb{E}\int_{Q}CzBz=2\mathbb{E}\sum_{i=1}^{6}\sum_{j=1}^{3}\int_{Q}C_{i}zB_{j}z$ and to bound the left-hand side of (B.2) with respect to the term $M_{h}z$. Indeed, We notice that from (B.2), it holds that
(B.3) $2r\mathcal{P}(q)\,Cz=2|Cz|^{2}dt+2Cz\,Bz-2M_{h}z\,Cz\,dt.$
Hence, combining the definition of $r\mathcal{P}(q)$ with (B.3), it follows that
$\mathbb{E}\int_{Q}2rf\,Cz\,dt=2\mathbb{E}\int_{Q}|Cz|^{2}dt+2\mathbb{E}\int_{Q}CzBz-2\mathbb{E}\int_{Q}M_{h}z\,Cz\,dt.$
Using the Young inequality, we notice that
(B.4) $2\mathbb{E}\int_{0}^{T}|rf|^{2}\,dt\geq\mathbb{E}\int_{Q}|Cz|^{2}dt+2\mathbb{E}\int_{Q}Cz\,Bz-2\mathbb{E}\int_{Q}|M_{h}z|^{2}\,dt.$
The next step is to provide an estimate for the right-hand side of (B.4).
Remark B.1.
The decomposition of (B.2) differs from that presented in EstimateCarlemanStochasticControl , not only due to the sign change, but also because the corrector, which is typically added when the dimension is greater than one, is included as an additional term rather than as a component of $M_{h}z$. As a result, the estimate for the cross-product includes an additional term that must be considered. Moreover, in the estimate (B.4), Young’s inequality is applied in a way that preserves the term $|C(z)|^{2}$, which will play a crucial role in the final estimate of the Carleman inequality. These are some of the reasons for the study of this new configuration in order to obtain the Carleman inequality.
B.2. An estimate for the cross-product
Recall that
(B.5) $2\mathbb{E}\int_{Q}CzBz=2\mathbb{E}\sum_{i=1}^{6}\sum_{j=1}^{3}\int_{Q}C_{i}zB_{j}z\triangleq\,\sum_{i=1}^{6}\sum_{j=1}^{3}I_{ij}.$
The procedure for estimating the product (B.5) is analogous to that in EstimateCarlemanStochasticControl , except for a change of sign and consideration of Remark B.1. Consequently, we divide the estimate into three groups: terms involving the differential $dz$, products of the additional terms, and terms involving the differential $dt$.
We begin with the terms involving the differential $dt$. Specifically, consider the terms $I_{i}$ for $i=1,2,3,\dots,5$, whose estimates can be found in EstimateCarlemanStochasticControl , up to a sign change. For the additional term $I_{6}$, which does not appear in EstimateCarlemanStochasticControl , its estimate is provided in C. Consequently, we obtain the following preliminary estimate:
Lemma B.2.
(Terms involving differential $dz$.) For $\max_{t\in[0,T]}\{\theta(t)\}\tau h\leq 1$, we have
$\begin{split}\sum_{i=1}^{6}I_{i1}\geq\,&-\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}\gamma_{i}|D_{i}(dz)|^{2}+\mathbb{E}\int_{Q}(s^{2}\lambda^{2}\varphi^{2}+s\lambda^{2}\varphi)|\nabla_{\gamma}\psi|^{2}|dz|^{2}-X_{1}-Y_{1}\end{split}$
where
$\begin{split}X_{1}\triangleq\,&\sum_{i=1}^{n}\left(\mathbb{E}\int_{Q_{i}^{\ast}}\mathcal{O}_{\lambda}((sh)^{2})\,|D_{i}(dz)|^{2}+\mathbb{E}\int_{Q_{i}^{\ast}}T\theta(t)\mathcal{O}_{\lambda}((sh)^{2})\,|D_{i}z|^{2}\,dt\right)\\ &+\mathbb{E}\int_{Q}(\mathcal{O}_{\lambda}((sh)^{2})+T\mathcal{O}_{\lambda}(1)+s\lambda\varphi\mathcal{O}(1))\,|dz|^{2}\\ &+\mathbb{E}\int_{Q}T(\theta\mathcal{O}_{\lambda}((sh)^{2})-2s\varphi(\lambda^{3}\beta|\nabla_{\gamma}\psi|^{2}+\lambda^{2}\mathcal{O}(1)))+\mathcal{O}_{\lambda}(1))\,|z|^{2}dt\end{split}$
and
$\begin{split}Y_{1}\triangleq\,&-\left.\mathbb{E}\int_{\mathcal{M}}[s^{2}\mathcal{O}_{\lambda}(1)+2s(\lambda^{2}\varphi|\nabla_{\gamma}\psi|^{2}+\lambda\varphi\mathcal{O}(1))]|z|^{2}\right|_{0}^{T}+\sum_{i=1}^{n}\left.\mathbb{E}\int_{\mathcal{M}_{i}}\gamma_{i}|D_{i}z|^{2}\right|_{0}^{T}\\ &\sum_{i=1}^{n}\mathbb{E}\left.\int_{\mathcal{M}_{i}^{*}}\mathcal{O}_{\lambda}((sh)^{2})\,|D_{i}z|^{2}\right|_{0}^{T}+\mathbb{E}\left.\int_{\mathcal{M}}(\mathcal{O}_{\lambda}((sh)^{2})+T\theta^{2}\mathcal{O}(1))\,|z|^{2}\right|_{0}^{T}.\end{split}$
Now, let us consider the products of the additional terms. In this case, the new terms $I_{62}$ and $I_{63}$ appear, whose estimates are provided in C. For the remaining terms, the estimates remain unchanged compared to EstimateCarlemanStochasticControl , since the weight functions still satisfy the same properties, namely $|\theta(t)|\leq C$ and $|\theta_{t}(t)|\leq C$ for some constant $C>0$ and all $t\in[0,T]$. Therefore, we obtain the following estimate:
Lemma B.3.
(product of the additional terms.) For $\max_{t\in[0,T]}\{\theta(t)\}\tau h\leq 1$, we obtain
$\sum_{i=4}^{6}I_{i2}+\sum_{i=1}^{6}I_{i3}\geq\,\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}4s\lambda^{2}\varphi\gamma_{i}|\nabla_{\gamma}\psi|^{2}\,|D_{i}z|^{2}\,dt-\mathbb{E}\int_{Q}4s^{3}\lambda^{4}\varphi^{3}|\nabla_{\gamma}\psi|^{4}\,|z|^{2}\,dt-X_{3}$
where
$X_{3}\triangleq\,\mathbb{E}\int_{Q}s|\mathcal{O}_{\lambda}((sh)^{2})|\,|z|^{2}\,dt-E\int_{Q}s\mathcal{O}_{\lambda}((sh)^{2})\,|z|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}s|\mathcal{O}_{\lambda}((sh)^{2})|\,|D_{i}z|^{2}\,dt$
$+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}s\mathcal{O}_{\lambda}((sh)^{2})\,|D_{i}z|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}(4s\lambda\Delta_{\gamma}\psi\,\varphi\gamma_{i}+s\mathcal{O}_{\lambda}(h+(sh)^{2}))\,|D_{i}z|^{2}\,dt$
$+\mathbb{E}\int_{Q}s\mathcal{O}_{\lambda}(1)\,|z|^{2}\,dt+\mathbb{E}\int_{Q}(s^{3}\lambda^{3}\varphi^{3}\mathcal{O}(1)+s^{2}\mathcal{O}_{\lambda}(1)+s^{3}\mathcal{O}_{\lambda}((sh)^{2}))\,|z|^{2}\,dt$
$+\mathbb{E}\int_{Q}(Cs\lambda^{2}T\theta^{2}\varphi|\nabla_{\gamma}\psi|^{2}+s\lambda T\theta^{2}\varphi\mathcal{O}(1))\,|z|^{2}\,dt{-2\mathbb{E}\int_{Q}s^{2}\lambda^{4}\varphi^{2}|\nabla_{\gamma}\psi|^{4}|z|^{2}\,dt}$
${+2\mathbb{E}\int_{Q}(s^{2}\lambda^{3}\varphi^{2}|\nabla_{\gamma}\psi|^{4}-4s^{2}\lambda^{2}|\mathcal{O}(1)|^{2}-s\mathcal{O}_{\lambda}(1)((sh)^{2}))\,|z|^{2}\,dt}.$
Finally, the terms $I_{12}$, $I_{22}$ and $I_{32}$ behave analogously to those in the deterministic setting studied in boyer-2014 , as the time variable does not play a significant role. For this reason, we omit a detailed proof of the following lemma, since the only difference with respect to boyer-2014 lies in our notation.
Lemma B.4.
(Terms involving differential $dt$.) For $\max_{t\in[0,T]}\{\theta(t)\}\tau h\leq 1$, we have the following inequality
$\sum_{i=1}^{3}I_{i2}\geq\,\mathbb{E}\int_{Q}6s^{3}\lambda^{4}\varphi^{3}|\nabla_{\gamma}\psi|^{4}\,|z|^{2}\,dt-\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}2s\lambda^{2}\varphi\gamma_{i}|\nabla_{\gamma}\psi|^{2}\,|D_{i}z|^{2}\,dt-X_{2}-Y_{2}$
where
$X_{2}\triangleq\,\sum_{i=1}^{n}\mathbb{E}\int_{Q}|s\lambda\varphi\mathcal{O}(1)+\mathcal{O}_{\lambda}(sh)+s\mathcal{O}_{\lambda}(sh)+s\mathcal{O}_{\lambda}((sh)^{2})|\,|D_{i}A_{i}z|^{2}\,dt$
$+\sum_{i=1}^{n}\mathbb{E}\int_{Q}h^{2}|\mathcal{O}_{\lambda}(sh)|\,|D_{i}^{2}z|^{2}\,dt+\sum_{i\neq j=1}^{n}\mathbb{E}\int_{Q_{ij}^{*}}\left|h\lambda\mathcal{O}(sh)+h\mathcal{O}_{\lambda}((sh)^{2})\right|\,|D_{ij}^{2}z|^{2}\,dt$
$+\sum_{i=1}^{n}\mathbb{E}\int_{Q^{\ast}_{i}}|s\lambda\varphi\mathcal{O}(1)+s\mathcal{O}_{\lambda}(sh)+h\mathcal{O}_{\lambda}(sh)+s\mathcal{O}_{\lambda}((sh)^{2})+\mathcal{O}_{\lambda}((sh)^{2})|\,|D_{i}z|^{2}$
$+\mathbb{E}\int_{Q}(s^{2}\lambda^{3}\varphi^{2}\mathcal{O}(1)+s^{2}\mathcal{O}_{\lambda}(1)+s^{2}T\theta\mathcal{O}_{\lambda}(1)+s^{3}\mathcal{O}_{\lambda}((sh)^{2}))|z|^{2}$
$+\sum_{i=1}^{n}\mathbb{E}\int_{Q}\left|h\lambda\mathcal{O}(sh)+h\mathcal{O}_{\lambda}((sh)^{2})\right|\,|D_{i}^{2}z|^{2}\,dt.$
and
$Y_{2}\triangleq\,$ $\sum_{i=1}^{n}\mathbb{E}\int_{\partial_{i}Q}\left(-2s\lambda\varphi(\gamma_{i})^{2}\partial_{i}\psi+s\mathcal{O}_{\lambda}(sh)+h\mathcal{O}_{\lambda}(sh)\right)t_{r}^{i}(|D_{i}z|^{2})\nu_{i}\,dt+$
$+\sum_{i=1}^{n}\mathbb{E}\int_{\partial_{i}Q}\mathcal{O}_{\lambda}(sh)t_{r}^{i}(|D_{i}z|^{2})\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{\partial_{i}Q}s\mathcal{O}_{\lambda}((sh)^{2})\,t_{r}^{i}(|D_{i}z|^{2})\nu_{i}\,dt$
Then by Lemma B.2-B.4 we obtain the following
(B.6) $\begin{split}2\mathbb{E}&\int_{Q}CzBz\geq\,\mathbb{E}\int_{Q}2s^{3}\lambda^{4}\varphi^{3}|\nabla_{\gamma}\psi|^{4}\,|z|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}2s\lambda^{2}\varphi\gamma_{i}|\nabla_{\gamma}\psi|^{2}\,|D_{i}z|^{2}\,dt\\ &-\sum_{i=1}^{n}\mathbb{E}\int_{Q^{\ast}}\gamma_{i}|D_{i}(dz)|^{2}+\mathbb{E}\int_{Q}(s^{2}\lambda^{2}\varphi^{2}+s\lambda^{2}\varphi)|\nabla_{\gamma}\psi|^{2}|dz|^{2}-\sum_{j=1}^{3}X_{j}-\sum_{j=1}^{2}Y_{j}.\end{split}$
To give an estimate of the right-hand side of (B.4), we need the following estimate $M_{h}z$. The proof can be adapted from Lemma 4.2 in boyer-2010-1d-elliptic and the estimation of $\Phi$ in zhao:2024 .
Lemma B.5.
(Estimate of $M_{h}z$.) For $\max_{t\in[0,T]}\{\theta(t)\}\tau h\leq 1$, we have
$\mathbb{E}\int_{Q}|M_{h}z|^{2}\,dt\leq\mathcal{O}_{\lambda}(1)\left(\mathbb{E}\int_{Q}s^{2}|z|^{2}\,dt+h^{2}\sum_{i=1}^{n}\int_{Q^{\ast}}s^{2}|D_{i}z|^{2}\,dt\right).$
In addition, we will need the following lemma, where its proof proceeds analogously to the proof of Lemma 3.11 in boyer-2014 .
Lemma B.6.
For $\max_{t\in[0,T]}\{\theta(t)\}\tau h\leq 1$, we have
$\begin{split}&\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}s\lambda^{2}\varphi\gamma_{i}|\nabla_{\gamma}\psi|^{2}|D_{i}z|^{2}\,dt\geq\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\int_{Q}s\lambda^{2}\varphi\gamma_{i}|\nabla_{\gamma}\psi|^{2}|A_{i}D_{i}z|^{2}\,dt+X_{3}+Y_{3},\end{split}$
where
$\begin{split}&X_{3}\triangleq\,\sum_{i\neq=1}^{n}\left(\mathbb{E}\int_{Q_{ij}^{\ast}}h\mathcal{O}_{\lambda}(sh)\,|D^{2}_{ij}z|^{2}\,dt-\mathbb{E}\int_{Q_{i}^{\ast}}h\mathcal{O}(sh)|D_{i}z|^{2}\,dt\right)\\ &+\sum_{i=1}^{n}\left(\mathbb{E}\int_{Q}h\mathcal{O}_{\lambda}(sh)|A_{i}D_{i}z|^{2}\,dt+\mathbb{E}\int_{Q}h\mathcal{O}_{\lambda}(sh)|D^{2}_{i}z|^{2}\,dt-\mathbb{E}\int_{Q_{i}^{\ast}}h\mathcal{O}_{\lambda}(sh)|D_{i}z|^{2}\,dt\right)\end{split}$
and
$Y_{3}\triangleq\,\sum_{i=1}^{n}\mathbb{E}\int_{\partial_{i}Q}h\mathcal{O}(sh)t_{r}^{i}(|D_{i}z|^{2})\,dt.$
Combining the lemma B.5 and B.6 with (B.6), we see that for $sh\leq\varepsilon$, there exist $\lambda_{1}\geq 1$, and $\varepsilon_{1}(\lambda)>0$ such that for $\lambda\geq\lambda_{1}$ and $0<sh\leq\varepsilon_{1}(\lambda)$, we have
[…]
with
(B.8) $\begin{split}\tilde{X}\triangleq&\sum_{l=1}^{2}X_{l}+Y_{l}-\mathbb{E}\int_{Q}s^{3}\lambda^{3}\varphi^{3}\mathcal{O}(1)|z|^{2}\,dt\end{split}-\mathbb{E}\int_{Q}s^{2}\lambda^{4}\varphi^{2}\mathcal{O}(1)\,|z|^{2}\,dt$
If we fix $\lambda=\lambda_{1}$, we can choose $\varepsilon_{0}$ and $h_{0}$ to be sufficiently small, with $0<\varepsilon\leq\varepsilon(\lambda_{1})$, and $s_{0}\geq 1$ sufficiently large, so that for $s\geq s_{0}$, $0<h\leq h_{0}$, and $sh\leq\varepsilon_{0}$, we obtain
[…]
where
(B.10) $\begin{split}\tilde{Y}\triangleq\,\sum_{i=1}^{n}\mathbb{E}\int_{\partial_{i}Q}2s\lambda_{1}\varphi(\gamma_{i})^{2}\partial_{i}\psi\,&t_{r}^{i}(|D_{i}z|^{2})\nu_{i}\,dt\\ &+C_{s_{0},\epsilon}\left(\left.-\sum_{i=1}^{n}\mathbb{E}\int_{\mathcal{M}_{i}^{\ast}}|D_{i}z|^{2}\right|_{0}^{T}+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,|z|^{2}\right|_{0}^{T}\right)\end{split}$
Moreover, since $z(0)=0$ and recalling that $\psi$ satisfies (2.1), we obtain the following
$\tilde{Y}=\,C_{s_{0},\epsilon}\left(\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}s\lambda t_{r}^{i}(|D_{i}z|^{2})\,dt-\left.\mathbb{E}\int_{\mathcal{M}_{i}^{\ast}}|D_{i}z|^{2}\right|_{t=T}+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}|z|^{2}\right|_{t=T}\right).$
Therefore, we write (B.9) as
(B.11) $\begin{split}&\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}|\nabla_{\gamma}\psi|^{4}\,|z|^{2}\,dt+\mathbb{E}\int_{Q}(s^{2}\lambda_{1}^{2}\varphi^{2}+s\lambda_{1}^{2}\varphi)|\nabla_{\gamma}\psi|^{2}|dz|^{2}+\mathbb{E}\int_{Q}|Cz|^{2}\,dt\\ &+\sum_{i=1}^{n}\left(\mathbb{E}\int_{Q_{i}^{\ast}}s\lambda^{2}_{1}|\nabla_{\gamma}\psi|^{2}\varphi\,|D_{i}z|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi|\nabla_{\gamma}\psi|^{2}|A_{i}D_{i}z|^{2}\,dt-\mathbb{E}\int_{Q_{i}^{\ast}}\gamma_{i}|D_{i}(dz)|^{2}\right)\\ &\leq\,C_{s_{0},\varepsilon}\left(\mathbb{E}\int_{Q}|rf|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}s\lambda_{1}t_{r}^{i}(|D_{i}z|^{2})\,dt+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}|z|^{2}\right|_{t=T}\right).\end{split}$
B.3. Returning to the original variable
We now return to our original variable. The methodology in this section departs from the approach in EstimateCarlemanStochasticControl . Therefore, as a first step, we establish the following result for the terms on the left-hand side of the desired Carleman estimate, which corresponds to the left-hand side of inequality (B.11).
Lemma B.7.
Let $\lambda=\lambda_{1}$, then there exists $\varepsilon_{2}$ and $h_{2}$ sufficiently small, with $0<\varepsilon\leq\varepsilon(\lambda_{1})$, and $s_{2}\geq 1$ sufficiently large, such that for $s\geq s_{2}$, $0<h\leq h_{2}$, and $sh\leq\varepsilon_{2}$, we obtain the following inequality:
(B.12) $\begin{split}J(w)+&\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}|rw|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda^{2}_{1}\varphi|rg|^{2}\,dt-\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}|rD_{i}(g)|^{2}\,dt\\ &\leq C^{\prime}_{s_{2},\epsilon_{2}}\left(\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}|z|^{2}\,dt+\mathbb{E}\int_{Q}(s^{2}\lambda^{2}\varphi^{2}+s\lambda^{2}\varphi)|\nabla_{\gamma}\psi|^{2}|dz|^{2}\right.\\ &\left.+\sum_{i=1}^{n}\left[\mathbb{E}\int_{Q}s\lambda^{2}\varphi|A_{i}D_{i}z|^{2}\,dt+\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi\lambda^{2}\,|D_{i}z|^{2}\,dt-\mathbb{E}\int_{Q_{i}^{\ast}}|D_{i}(dz)|^{2}\right]\right)\end{split}$
where $J(w)\triangleq\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi\lambda^{2}|rD_{i}w|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda^{2}\varphi|rA_{i}D_{i}w|^{2}\,dt$ and $w$ is solution of $dw+\sum_{i=1}^{n}D_{i}(\gamma_{i}w)dt=fdt+gdB(t)$ with $w=0$ on $\partial Q$.
Proof.
Recalling $w=\rho z$, we have $D_{i}w=A_{i}\rho\,D_{i}z+D_{i}\rho\,A_{i}z$. Therefore, by Lemma A.1 and A.3, for each $i\in\llbracket 1,n\rrbracket$ it holds that
(B.13) $\begin{split}\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi\lambda^{2}|rD_{i}w|^{2}\,dt\leq\;&\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi\lambda^{2}\,|D_{i}z|^{2}\,dt+\mathbb{E}\int_{Q_{i}^{\ast}}s\mathcal{O}_{\lambda}((sh)^{2})\,|D_{i}z|^{2}\,dt\\ &\mathbb{E}\int_{Q_{i}^{\ast}}s^{3}\varphi^{3}\lambda^{4}|A_{i}z|^{2}\,dt+\mathbb{E}\int_{Q_{i}^{\ast}}s^{3}\mathcal{O}_{\lambda}((sh)^{2})\,|A_{i}z|^{2}\,dt.\end{split}$
Now, using (2.5), and by the integration by parts with respect average operator, we have
$\begin{split}\mathbb{E}\int_{Q_{i}^{\ast}}s^{3}\varphi^{3}\lambda^{4}|A_{i}z|^{2}\,dt\leq\mathbb{E}\int_{Q_{i}^{\ast}}s^{3}\varphi^{3}\lambda^{4}A_{i}(|z|^{2})\,dt=\mathbb{E}\int_{Q}s^{3}A_{i}(\varphi^{3})\lambda^{4}|z|^{2}\,dt,\end{split}$
where used that $z=0$ on $\partial_{i}Q$. Applying the approximation to function $A_{i}(\varphi^{3})=\varphi^{3}+\mathcal{O}_{\lambda}((sh)^{2})$ on integral above, we obtain that
(B.14) $\mathbb{E}\int_{Q_{i}^{\ast}}s^{3}\varphi^{3}\lambda^{4}|A_{i}z|^{2}\,dt\leq\mathbb{E}\int_{Q}s^{3}\varphi^{3}\lambda^{4}|z|^{2}\,dt+\mathbb{E}\int_{Q}s\mathcal{O}_{\lambda}((sh)^{2})|z|^{2}\,dt.$
Similar to the derivation above, we see that
(B.15) $\mathbb{E}\int_{Q_{i}^{\ast}}s^{3}\mathcal{O}_{\lambda}((sh)^{2})\,|A_{i}z|^{2}\,dt\leq\mathbb{E}\int_{Q}s^{3}\mathcal{O}_{\lambda}((sh)^{2})\,|z|^{2}\,dt.$
Then, combining (B.13)-(B.15), we obtain the following inequality for first term of the $J(w)$:
(B.16) $\begin{split}\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi\lambda^{2}&|rD_{i}w|^{2}\,dt\leq\;\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi\lambda^{2}\,|D_{i}z|^{2}\,dt+\mathbb{E}\int_{Q_{i}^{\ast}}s\mathcal{O}_{\lambda}((sh)^{2})\,|D_{i}z|^{2}\,dt\\ &\mathbb{E}\int_{Q}s^{3}\varphi^{3}\lambda^{4}|z|^{2}\,dt+\mathbb{E}\int_{Q}s\mathcal{O}_{\lambda}((sh)^{2})|z|^{2}\,dt+\mathbb{E}\int_{Q}s^{3}\mathcal{O}_{\lambda}((sh)^{2})\,|z|^{2}\,dt.\end{split}$
Now, for the second terms of the $J(w)$ we use (1.13) and (1.14), therefore
$\begin{split}A_{i}(D_{i}(w))=&A_{i}(D_{i}(\rho)A_{i}(z))+A_{i}(A_{i}(\rho)D_{i}(z))\\ =&A_{i}D_{i}(\rho)\,A_{i}^{2}(z)+A_{i}^{2}(\rho)\,A_{i}D_{i}(z)+\frac{h^{2}}{4}\left(D_{i}^{2}(\rho)\,D_{i}A_{i}(z)+D_{i}A_{i}(\rho)\,D_{i}^{2}(z)\right).\end{split}$
Applying (1.15) to the first term on the right side of the previous equation, we obtain the following equality
(B.17) $rA_{i}D_{i}w=rA_{i}D_{i}(\rho)\,z+rA_{i}^{2}(\rho)\,A_{i}D_{i}(z)+\frac{h^{2}}{4}\left(rD_{i}^{2}(\rho)\,D_{i}A_{i}(z)+2rD_{i}A_{i}(\rho)\,D_{i}^{2}(z)\right).$
With Proposition A.3 and Lemma A.1 we then find
(B.18) $rA_{i}D_{i}w=A_{i}D_{i}z-s\lambda\varphi\mathcal{O}(1)\,z+\mathcal{O}_{\lambda}((sh)^{2})\,D_{i}A_{i}z+h\mathcal{O}_{\lambda}(sh)\,D_{i}^{2}(z)$
Now, by applying the triangle inequality and (B.18), we obtain the following estimate for each $i\in\llbracket 1,d\rrbracket$ of the second terms from $J(w)$:
(B.19) $\begin{split}\mathbb{E}\int_{Q}s\lambda^{2}\varphi&|rA_{i}D_{i}w|^{2}\,dt\\ \leq&\,\mathbb{E}\int_{Q}s\lambda^{2}\varphi|A_{i}D_{i}z|^{2}\,dt+\mathbb{E}\int_{Q}s^{3}\lambda^{4}\varphi\mathcal{O}(1)\,|z|^{2}\,dt+\mathbb{E}\int_{Q}\mathcal{O}_{\lambda}((sh)^{4})\,|D^{2}_{i}z|^{2}\,dt.\end{split}$
Combining (B.16) and (B.19), we can estimate to $J(w)$ as
(B.20) $\begin{split}J(w)\leq\sum_{i=1}^{n}&\left[\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi\lambda^{2}\,|D_{i}w|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda^{2}\varphi|A_{i}D_{i}z|^{2}\,dt+\mathbb{E}\int_{Q_{i}^{\ast}}s\mathcal{O}_{\lambda}((sh)^{2})\,|D_{i}z|^{2}\,dt\right.\\ &\left.+\mathbb{E}\int_{Q}\mathcal{O}_{\lambda}((sh)^{4})\,|D^{2}_{i}z|^{2}\,dt\right]+\mathbb{E}\int_{Q}s^{3}\lambda^{4}\varphi\mathcal{O}(1)\,|z|^{2}\,dt.\end{split}$
Now, we will focus on estimating the two terms on the left-hand side of the inequality in this Lemma. Given the earlier definition $w=\rho z$ we can derive the relationship between their differentials: $dw=\rho\,dz+\partial_{t}(\rho)\,zdt$. Consequently, applying Proposition 1.8 we obtain
$rA_{i}(\rho)D_{i}(dz)=rD_{i}(dw)-rD_{i}(\partial_{t}(\rho))A_{i}(z)dt-rA_{i}(\partial_{t}(\rho))D_{i}(z)dt-rD_{i}(\rho)A_{i}(dz).$
Now, by (A.4), (A.1) and recall that $r\rho=1$ it follows that
$D_{i}(dz)=$ $rD_{i}(dw)-\mathcal{O}_{\lambda}((sh)^{2})D_{i}(dz)+s^{2}\lambda^{2}\varphi^{2}\theta_{t}(\partial_{i}\psi)A_{i}(z)dt-s\lambda^{2}\varphi\theta_{t}(\partial_{i}\psi)A_{i}(z)dt$
$-s\lambda\varphi\theta_{t}D_{i}(z)dt+s\varphi\lambda(\partial_{i}\psi)A_{i}(dz)+\mathcal{O}_{\lambda}((sh)^{2})D_{i}z\,dt-s\mathcal{O}_{\lambda}((sh)^{2})A_{i}(dz)$
Therefore, from the above equation for each $i\in\llbracket 1,n\rrbracket$, we obtain that
$\begin{split}\mathbb{E}\int_{Q_{i}^{\ast}}\gamma_{i}|D_{i}(dz)|^{2}=&\mathbb{E}\int_{Q_{i}^{\ast}}\gamma_{i}|rD_{i}(w)|^{2}+\mathbb{E}\int_{Q_{i}^{\ast}}\mathcal{O}_{\lambda}((sh)^{4})|D_{i}(dz)|^{2}\\ &+\mathbb{E}\int_{Q_{i}^{\ast}}s^{2}\lambda^{2}\gamma_{i}(\partial_{i}\psi)^{2}|A_{i}(dz)|^{2}+\mathbb{E}\int_{Q_{i}^{\ast}}s^{2}\mathcal{O}_{\lambda}((sh)^{4})|A_{i}(dz)|^{2}.\end{split}$
Similar to (B.14) and (B.15), we find that
(B.21) $\begin{split}\mathbb{E}\int_{Q_{i}^{\ast}}\gamma_{i}|D_{i}(dz)|^{2}\leq&\mathbb{E}\int_{Q_{i}^{\ast}}\gamma_{i}|rD_{i}(w)|^{2}+\mathbb{E}\int_{Q}s^{2}\lambda^{2}\varphi^{2}(\partial_{i}\psi)^{2}|dz|^{2}\\ &+\mathbb{E}\int_{Q_{i}^{\ast}}\mathcal{O}_{\lambda}((sh)^{4})|D_{i}(dz)|^{2}+\mathbb{E}\int_{Q}\mathcal{O}_{\lambda}(sh)^{4})|dz|^{2}.\end{split}$
Noting that
$\mathbb{E}\int_{Q}s\lambda^{2}\varphi\gamma_{i}(\partial_{i}\psi)^{2}|dz|^{2}=\mathbb{E}\int_{Q}s\lambda^{2}\varphi\gamma_{i}(\partial_{i}\psi)^{2}|rdw|^{2}$
Then, from (B.21) we obtain for each $i\in\llbracket 1,n\rrbracket$
(B.22) $\begin{split}&\mathbb{E}\int_{Q}s\lambda^{2}\varphi\gamma_{i}(\partial_{i}\psi)^{2}|rdw|^{2}-\mathbb{E}\int_{Q_{i}^{\ast}}\gamma_{i}|rD_{i}(w)|^{2}\leq-\mathbb{E}\int_{Q_{i}^{\ast}}\gamma_{i}|D_{i}(dz)|^{2}\\ &+\mathbb{E}\int_{Q_{i}^{\ast}}\mathcal{O}_{\lambda}((sh)^{4})|D_{i}(dz)|^{2}+\mathbb{E}\int_{Q}\left(s^{2}\lambda^{2}\varphi^{2}+s\lambda^{2}\varphi\right)\gamma_{i}(\partial_{i}\psi)^{2}|dz|^{2}+\mathbb{E}\int_{Q}\mathcal{O}_{\lambda}(sh)^{4})|dz|^{2}.\end{split}$
Du to the inequalities (B.20),(B.28) and (B.22), we conclude the proof of Lemma (B.7).This shows that for $0<sh\leq\varepsilon$ and $\lambda=\lambda_{1}$, there exist $\varepsilon_{2}(\lambda)>0$ and $s_{2}\geq 1$ such that for $s\geq s_{2}$, $0<h<h_{2},$ and $0<sh\leq\varepsilon_{2}(\lambda)$, the desired inequality is obtained. ∎
To proceed, we introduce the following lemma, which allows us to incorporate the second-order spatial term $D_{ij}^{2}$ into the left-hand side of the Carleman estimate. The proof follows by adapting the approach in (lecaros2025lipschitzstabilityinverseproblems, , Theorem 1.2).
Lemma B.8.
There exists $\varepsilon_{3}(\lambda)>0$ and $s_{3}\geq 1$ such that for $s\geq s_{3}$, $0<h<h_{3}$, and $0<sh\leq\varepsilon_{3}(\lambda)$, we obtain the following inequality:
$\begin{split}&\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}w|^{2}\,dt-\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}}s\lambda^{2}\varphi^{2}|\nabla\psi|_{\gamma}^{2}\gamma_{i}\partial_{i}\psi\,|rD_{i}w|^{2}+\mathcal{O}(s^{-1})|rA_{i}D_{i}w|^{2}\,dt\\ &\leq C_{\varepsilon_{3},s_{3}}\left(\mathbb{E}\int_{Q}|Cz|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda^{4}\varphi^{2}|\nabla_{\gamma}\psi|^{2}|z|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q}s\lambda^{2}\varphi^{2}|\nabla_{\gamma}\psi|^{2}\,|A_{i}D_{i}z|^{2}\,dt\right)\end{split}$
Proof.
Recalling that $r\sum_{i\in\llbracket 1,n\rrbracket}D_{i}(\gamma_{i}D_{i}w)=Cz+B_{2}z+B_{3}z-M_{h}z\triangleq\mathcal{B}w$ and applying Young’s inequality, we obtain
(B.23) $|\mathcal{B}w|^{2}\leq 2(|Cz|^{2}+|B_{2}z|^{2}+|B_{3}z|^{2}+|M_{h}z|^{2})$
Therefore, taking $s>1$ we have a first estimate for the third left-side term of (B.12) given by
(B.24) $\begin{split}\mathbb{E}\int_{Q}s^{-1}|\mathcal{B}w|^{2}\,dt\leq\mathbb{E}\int_{Q}2|Cz|^{2}\,dt+&\mathbb{E}\int_{Q}2s^{-1}|B_{2}z|^{2}\,dt\\ &+\mathbb{E}\int_{Q}2s^{-1}|B_{3}z|^{2}\,dt+\mathbb{E}\int_{Q}2s^{-1}|M_{h}z|^{2}\,dt.\end{split}$
From the definition of the $B_{2}z$ and $B_{3}z$ respectively, Lemma A.1 and Proposition A.3, we can see that
(B.25) $\begin{split}\mathbb{E}\int_{Q}2s^{-1}|B_{2}z|^{2}dt\leq&\sum_{i=1}^{n}\mathbb{E}\int_{Q}8s^{-1}|\gamma_{i}rD_{i}A_{i}\rho\,D_{i}A_{i}z|^{2}dt\\ \leq&8\sum_{i=1}^{n}\mathbb{E}\int_{Q}(s\lambda^{2}\varphi^{2}|\nabla_{\gamma}\psi|^{2}+\mathcal{O}_{\lambda}|((sh)^{4})||A_{i}D_{i}z|^{2}dt,\end{split}$
(B.26) $\mathbb{E}\int_{Q}2s^{-1}|B_{3}z|^{2}dt=4\mathbb{E}\int_{Q}2s^{-1}|s\Delta_{\gamma_{i}}\varphi\,z|^{2}dt\leq 8\mathbb{E}\int_{Q}(s\lambda^{4}\varphi^{2}|\nabla_{\gamma}\psi|^{4}+s\lambda|\mathcal{O}((1)|^{2}|z|^{2}\;dt$
and similar to Lemma B.5, we can deduce the following estimate for $M_{h}z$ as
(B.27) $\mathbb{E}\int_{Q}2s^{-1}|M_{h}z|^{2}\,dt\leq\mathcal{O}_{\lambda}(1)\left(\mathbb{E}\int_{Q}s|z|^{2}\,dt+h^{2}\sum_{i=1}^{n}\int_{Q^{\ast}}s|D_{i}z|^{2}\,dt\right).$
Thus, combining (B.24)-(B.27) we obtain the estimate
(B.28) $\begin{split}\mathbb{E}\int_{Q}&s^{-1}|\mathcal{B}w|^{2}\,dt\\ \leq&\;\mathbb{E}\int_{Q}|Cz|^{2}dt+\mathbb{E}\int_{Q}(4s\lambda^{4}\varphi^{2}|\nabla_{\gamma}\psi|^{4}+s\mathcal{O}_{\lambda}(1)+s\lambda|\mathcal{O}((1)|^{2}|z|^{2}dt\\ &+\sum_{i=1}^{n}\left(\mathbb{E}\int_{Q}h\mathcal{O}_{\lambda}(sh)|D_{i}z|^{2}dt+\mathbb{E}\int_{Q}(4s\lambda^{2}\varphi^{2}|\nabla_{\gamma}\psi|^{2}+\mathcal{O}_{\lambda}|((sh)^{4})||A_{i}D_{i}z|^{2}dt\right).\end{split}$
In turn, our next task is to compare the terms $\mathbb{E}\int_{Q}s^{-1}|\mathcal{B}w|^{2}\,dt$ and
$\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{*}}s^{-1}\gamma_{i}\gamma_{j}r^{2}|D_{ij}w|^{2}\;dt$. To this end, we can now proceed analogously to the proof of Theorem 1.2 in lecaros2025lipschitzstabilityinverseproblems . In particular, from the inequalities (2.27)-(2.32) in lecaros2025lipschitzstabilityinverseproblems it follows that
(B.29) $\begin{split}\mathbb{E}\int_{Q}s^{-1}|\mathcal{B}w|^{2}\,dt\geq&\sum_{i,j\in\llbracket 1,n\rrbracket}\int_{Q_{ij}^{\ast}}s^{-1}\gamma_{i}\gamma_{j}|rD_{ij}^{2}w|^{2}\,dt-\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}}s\lambda^{2}\varphi^{2}|\nabla\psi|_{\gamma}^{2}\gamma_{i}\partial_{i}\psi|rD_{i}w|^{2}\,dt\\ &-\sum_{i=1}^{n}\mathbb{E}\int_{Q}s^{-1}r^{2}|\mathcal{O}(1)|^{2}|A_{i}D_{i}w|^{2}\,dt-K(w),\end{split}$
where $K(w)$ is given by
$K(w)\triangleq$ $\sum_{i,j\in\llbracket 1,n\rrbracket}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1/2}\left(\mathcal{O}(h)+\mathcal{O}_{\lambda}((sh)^{2}+\mathcal{O}_{\lambda}(1)\right)|rD_{ij}^{2}w|^{2}\,dt$
$+\sum_{j\in\llbracket 1,n\rrbracket}\mathbb{E}\int_{Q_{j}^{\ast}}\left(s^{-1/2}\left(\mathcal{O}(h)+\mathcal{O}_{\lambda}((sh)^{2}+\mathcal{O}_{\lambda}(1)\right)+\mathcal{O}_{\lambda}(1)+s\mathcal{O}(sh)\right)|rD_{j}w|^{2}\,dt.$
Finally, there exist $\varepsilon_{3}(\lambda)>0$ and $s_{3}\geq 1$ such that combining (B.29) and (B.28) for $s\geq s_{3}$, $0<h<h_{3}$, and $0<sh\leq\varepsilon_{3}(\lambda)$, the desired inequality is obtained ∎
The two previous lemmas allow us to obtain a lower bound for the inequality (B.11) in terms of $w$. Now, we need to obtain an upper bound for the same inequality. To achieve this, we only need to estimate the second term on the right-hand side of the inequality (B.11), since the other terms are immediately available in terms of $w$.
Recalling that $w=\rho w$ and by Proposition 1.8 and Lemma A.1, we have $rD_{i}w=D_{i}z+\mathcal{O}_{\lambda}((sh)^{2})D_{i}z-\lambda s\varphi(\partial_{i}\psi)A_{i}(z)+s\mathcal{O}_{\lambda}((sh)^{2})A_{i}(z)$. We obtain the following
$\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}$ $s\lambda_{1}\varphi t_{r}^{i}(|D_{i}z|^{2})\,dt\leq\sum_{i=1}^{n}\left(\mathbb{E}\int_{0}^{T}\int_{\Gamma_{+}}s\lambda_{1}\varphi t_{r}^{i}(|rD_{i}(w)|^{2})dt\right$.
$\left.+\mathbb{E}\int_{0}^{T}\int_{\Gamma_{+}}s\mathcal{O}_{\lambda}((sh)^{2})t_{r}^{i}(D_{i}z)dt+\mathbb{E}\int_{0}^{T}\int_{\Gamma_{+}}\lambda^{3}s^{3}\varphi^{3}(\partial_{i}\psi)^{2}t_{r}^{i}(|A_{i}(z)|^{2})dt\right$.
$\left.+\mathbb{E}\int_{0}^{T}\int_{\Gamma_{+}}s^{3}|\mathcal{O}_{\lambda}((sh)^{2})|^{2}t_{r}^{i}(|A_{i}(z)|^{2})dt\right).$
Now, as $z=0$ on $\partial Q$ we notice that $t_{r}^{i}(|A_{i}z|^{2})=\frac{h^{2}}{2}t_{r}^{i}(|D_{i}z|^{2})$. Then, we can see that
$\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}$ $\int_{\Gamma^{+}}s\lambda_{1}\varphi t_{r}^{i}(|D_{i}z|^{2})\,dt\leq\sum_{i=1}^{n}\left(\mathbb{E}\int_{0}^{T}\int_{\Gamma_{+}}s\lambda_{1}\varphi t_{r}^{i}(|rD_{i}(w)|^{2})dt\right$.
$\left.+\mathbb{E}\int_{0}^{T}\int_{\Gamma_{+}}s\mathcal{O}_{\lambda}((sh)^{2})t_{r}^{i}(D_{i}z)dt+\mathbb{E}\int_{0}^{T}\int_{\Gamma_{+}}s|\mathcal{O}_{\lambda}((sh)^{6})|t_{r}^{i}(|D_{i}(z)|^{2})dt\right).$
Therefore, by the previous inequality we can find $\varepsilon_{3}(\lambda)>0$ and $s_{3}>1$ such that for $s\geq s_{3}$, $0<h<h_{3}$ and $0<sh\leq\varepsilon_{3}(\lambda)$, we have
$\mathbb{E}\int_{Q}$ $|rf|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}s\lambda_{1}t_{r}^{i}(|D_{i}z|^{2})\,dt+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}|z|^{2}\right|_{t=T}$
$\leq$ $C^{\prime\prime}_{s_{3},\epsilon_{3}}\left(\mathbb{E}\int_{Q}|rf|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma_{+}}s\lambda_{1}\varphi t_{r}^{i}(|rD_{i}(w)|^{2})dt+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}|rw|^{2}\right|_{t=T}\right)$
Consequently, applying the inequality mentioned above, (B.11), Lemma B.7 and Lemma B.8, we can determine that for $s_{0}=\max\{s_{1},s_{2},s_{3}\}$ and $h_{0}=\min\{h_{1},h_{2},h_{3}\}$, the desired Carleman estimate is obtained.
Appendix C Technical steps for the cross-product estimate
This section derives additional estimates not covered in EstimateCarlemanStochasticControl , which are essential to establish lemmas B.2 and B.3.
We begin by estimating $I_{61}$. Let us define $C_{6}z=-2s\Delta_{\gamma}(\varphi)z$ and $B_{1}z=dz$. Then
$2\mathbb{E}\int_{Q}C_{6}zB_{1}z=-2\mathbb{E}\int_{Q}2s\Delta_{\gamma}(\varphi)zdz\triangleq\,I_{61}.$
Applying Itô’s formula yields the following result
$I_{61}=-2\left.\mathbb{E}\int_{Q}s\Delta_{\gamma}(\varphi)|z|^{2}\right|_{0}^{T}+2\mathbb{E}\int_{Q}\partial_{t}(s)\Delta_{\gamma}(\varphi)|z|^{2}\,dt+\mathbb{E}\int_{Q}s\Delta(\varphi)|dz|^{2}.$
Observing that $\Delta_{\gamma}(\varphi)=\lambda^{2}\varphi|\nabla_{\gamma}\psi|^{2}+\lambda\varphi\mathcal{O}(1)$ and $|\theta_{t}|\leq\lambda\beta T\theta$, we deduce
$I_{61}\geq\,\mathbb{E}\int_{Q}s\lambda^{2}\varphi|\nabla_{\gamma}\psi|^{2}|dz|^{2}+X_{61}$
with
$X_{61}\triangleq$ $-2\left.\mathbb{E}\int_{Q}s(\lambda^{2}\varphi|\nabla_{\gamma}\psi|^{2}+\lambda\varphi\mathcal{O}(1))|z|^{2}\right|_{0}^{T}+\mathbb{E}\int_{Q}s\lambda\varphi\mathcal{O}(1)|dz|^{2}$
$-2\mathbb{E}\int_{Q}s^{2}\lambda\beta T(\lambda^{2}\varphi|\nabla_{\gamma}\psi|^{2}+\lambda\varphi\mathcal{O}(1))|z|^{2}\,dt.$
We now proceed to estimate $I_{62}$. For this purpose, we set $C_{6}z=-2s\Delta_{\gamma}(\varphi)z$, $B_{2}z\,dt=-2\sum_{i=1}^{n}\gamma_{i}rD_{i}A_{i}\rho\,D_{i}A_{i}z\,dt.$ Defining $\beta_{62}^{i}=4s\gamma_{i}\Delta_{\gamma}(\varphi)rD_{i}A_{i}\rho$ and applying the discrete integration by parts formula (1.11), we obtain the following
$\mathbb{E}\int_{Q}\beta_{62}^{i}\,z\,D_{i}(A_{i}z)\,dt$ $=\mathbb{E}\int_{Q_{i}^{\ast}}D_{i}(\beta_{62}^{i}z)A_{i}z\,dt+\int_{\partial_{i}Q}\beta_{62}^{i}t_{r}^{i}(A_{i}z)z\,\nu_{i}\,dt$
$=\mathbb{E}\int_{Q_{i}^{\ast}}D_{i}(\beta_{62}^{i})|A_{i}z|^{2}\,dt-\frac{1}{2}\mathbb{E}\int_{Q_{i}^{\ast}}A_{i}(\beta_{62}^{i})D_{i}(|z|^{2})\,dt,$
where we used $z=0$ on $\partial_{i}Q$ and the product rule (1.13).
Next, applying the inequality (2.5) and the integration by parts formula for the average operator (1.12), we arrive at
$\mathbb{E}\int_{Q}\beta_{62}^{i}\,z\,D_{i}(A_{i}z)\,dt$ $\geq\mathbb{E}\int_{Q_{i}^{\ast}}D_{i}(\beta_{62}^{i})A(|z|^{2})\,dt+\frac{1}{2}\int_{Q}D_{i}A_{i}(\beta_{62}^{i})|z|^{2}\,dt$
$\geq-\frac{1}{2}\mathbb{E}\int_{Q}A_{i}D_{i}(\beta_{62}^{i})|z|^{2}\,dt-\frac{h}{2}\int_{\partial_{i}Q}|z|^{2}\,t_{r}^{i}(D_{i}(\beta_{62}^{i}))\,dt.$
Using the results of Appendix A and the boundary condition $z=0$ on $\partial_{i}Q$, we finally deduce
$I_{62}\geq 2\,\mathbb{E}\int_{Q}\big{(}s^{2}\lambda^{3}\varphi^{2}+s^{2}\lambda^{4}\varphi^{2}\big{)}|\nabla_{\gamma_{i}}\psi|^{4}|z|^{2}\,dt+\mathbb{E}\int_{Q}s\,\mathcal{O}_{\lambda}(1)(sh)^{2}.$
Finally, we focus on the estimate of $I_{63}$. Recall that $B_{3}z=-C_{6}z=-2s\Delta_{\gamma}(\varphi)z.$ Using the expression $\Delta_{\gamma}(\varphi)=\lambda^{2}\varphi|\nabla_{\gamma}\psi|^{2}+\mathcal{O}_{\lambda}(1)$ and the triangle inequality, we obtain the following.
$I_{63}=-4\,\mathbb{E}\int_{Q}s^{2}|\Delta_{\gamma}\varphi|^{2}\,|z|^{2}\,dt\geq-4\,\mathbb{E}\int_{Q}s^{2}\lambda^{4}\varphi^{2}|\nabla_{\gamma}\psi|^{4}\,|z|^{2}\,dt-4\,\mathbb{E}\int_{Q}s^{2}|\mathcal{O}_{\lambda}(1)|^{2}\,|z|^{2}\,dt.$